### v1

# Anisotropy can make a moving active fluid membrane rough or crumpled

Abstract
We present a hydrodynamic theory of anisotropic and inversion-asymmetric moving active permeable fluid membranes. These are described by an anisotropic Kardar-Parisi-Zhang equation. Depending upon the anisotropy parameters, the membrane can be large-scale anisotropic and logarithmically rough with translational quasi long range order and orientational long range order, together with the relaxational dynamics being logarithmically faster than ordinary diffusion. For other choices of the anisotropy parameters, the membrane is either effectively isotropic and algebraically rough with translational short, but orientational long range order, or crumpled.
The generality of the observations about the dynamical properties of the plasma membranes in live biological cells Alberts et al. (1994); Ramaswamy (2010); Joanny and Prost (2009); Marchetti et al. (2013); Levin and Korenstein (1991); Tuvia et al. (1997) has prompted scientists to construct minimal physical descriptions. It is, for instance, now understood that the macroscopic membrane properties in mixed lipid membranes depend on whether the membrane is in equilibrium Banerjee and Basu (2015), or active Banerjee and Basu (2018). In this Letter, we set up the hydrodynamic theory of a nearly flat patch of an anisotropic and inversion-asymmetric moving permeable active fluid membrane. The activity could be due to an active species living on the membrane, or due to activity in the embedding medium. For simplicity, we assume that there are no conservation laws or broken symmetry modes associated with the activity. Thus, if the activity is due to an active species, e.g., a membrane-spanning active protein, its concentration is not assumed to be conserved, or if it is due to the activity of the embedding medium, e.g., an orientable active fluid, it is in its isotropic phase. We consider an anisotropic membrane. For simplicity, we assume a fixed background for the membrane, i.e., the embedding fluid momentum is not conserved. We argue that the hydrodynamic equation for the undulation modes of such a membrane given by a single-valued height field $h({\bf x},t)$ is given by
$\frac{\partial h}{\partial t}+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (1)
Both $\lambda_{ij}$ and $\nu_{ij}$ are symmetric matrices. Noise $f$ has zero-mean and a variance given by
$\langle f({\bf x},t)f(0,0)\rangle=2D\delta^{2}(x)\delta(t).$ (2)
In the absence of any detailed balance, the noise strength $D$ has no specific relation to the damping coefficient matrix $\nu_{ij}$. For arbitrary $\nu_{ij}$ and $\lambda_{ij}$, Eq. (11) is an anisotropic generalization of the well-known isotropic Kardar-Parisi-Zhang (KPZ) equation Barabási and Stanley (1995). For simplicity and to reduce the number of parameters, we further impose invariance of Eq. (11) separately under $(x,y)\rightarrow(\pm y,\pm x)$ in two dimensions (2D). We thus set $\nu_{xx}=\nu_{yy}=\nu>0$, $\nu_{xy}=\nu_{yx}=\overline{\nu}$, $\lambda_{xx}=\lambda_{yy}=\lambda$ and $\lambda_{xy}=\lambda_{yx}=\overline{\lambda}$, making it distinct from the one considered earlier Wolf (1991); Täuber and Frey (2002); Altman et al. (2015).
The most prominent result from direct numerical simulations (DNS) of Eq. (11) in two dimensions (2D) is that it admits membrane conformations $h({\bf x},t)$ having strikingly different statistical properties controlled by the model parameters. See Figs. 1(a) and (b) and Figs. 2(a) and (b) for representative snapshots of $h({\bf x},t)$ falling in three distinct phases (color bars represent $h$ relative to its mean) depending on the degree of the surface anisotropy controlled by $\overline{\nu},\overline{\lambda}$; see also the corresponding movies MOV1 [for Fig. 1(a)], MOV2 [for Fig. 1(b)], MOV3(a) [for Fig. 2(a)] and MOV3(b)[for Fig. 2(b)] in Supplemental Material (SM) sup . We characterize these phases by numerically solving Eq. (11) and measuring width ${\mathcal{W}}\equiv\sqrt{\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle}$, which as a function of time $t$ has an initial growing part, followed by a saturated part ${\cal W}_{\text{sat}}$ that fluctuates about a steady mean Barabási and Stanley (1995); $\overline{h}(t)$ is the mean height at time $t$. We find ${\mathcal{W}}_{\text{sat}}$ to scale as $(\ln L/a_{0})^{\tilde{\chi}}$ and $L^{\chi}$, respectively, for Fig. 1(a) and Fig. 1(b); $\tilde{\chi}\approx 0.43\pm 0.019,\chi=0.379\pm 0.008$ in our DNS studies, $L$ and $a_{0}$ are system size and small scale cutoff, respectively. See Fig. 3. Hence we name them logarithmic rough and algebraic rough phases respectively. For Fig. 2(a) and (b), no steady states are reached (within the computation time), giving ever increasing fluctuations, suggesting a complete loss of order, and hence we call it the crumpled phase. Snapshots in Figs. 1(c) and (d) for respectively the logarithmic and algebraic rough phases for different sets of model parameters over enhanced ranges as used in Figs. 2(a) and (b) for the crumpled phase indicate significantly larger fluctuations in the latter. See later for more details on DNS.
Figure 1: Snapshots of the height profiles ($L=200$) in the steady states for (a) and (c) logarithmic rough phase, with parameters (a) $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$, and $\overline{\lambda}=4$, and (c) $\nu=0.60$, $\overline{\nu}=0.35$, $\lambda=5$, and $\overline{\lambda}=8$. Snapshots (b) and (d) correspond to the algebraic rough phase, with parameters (b) $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$, and $\overline{\lambda}=1$, and (d) $\nu=0.60$, $\overline{\nu}=0.35$, $\lambda=20$, and $\overline{\lambda}=2$. Notice the much larger fluctuations in the algebraic rough phase. See text.
Figure 2: Snapshots of $h({\bf x},t)$ ($L=200$) in the crumpled phase with $\nu=1,\lambda=2$ showing much larger fluctuations than those in Fig. 1. (a) $\overline{\nu}=2,\overline{\lambda}=2$ (with ripples at $\pi/4$), (b) $\overline{\nu}=-2,\overline{\lambda}=-2$ (with ripples at $43\pi/$ 4). See text.
We show that the logarithmic rough phase is characterized by translational quasi-long-range order (QLRO) and orientational long-range order (LRO) with relaxation time scales being logarithmically faster than ordinary diffusion. Such a surface is effectively inversion-symmetric but has large-scale anisotropy. Further, the algebraically rough phase has effective isotropy with inversion-asymmetry and translational SRO and orientational LRO, akin to a 2D KPZ surface Kardar et al. (1986); Tang et al. (1990); Barabási and Stanley (1995) and crumpled phase with translational and orientational SRO, respectively. These results should be contrasted with the corresponding isotropic case, which only has a rough phase.
We now derive Eq. (11) and use it to establish the above results and obtain the phase diagrams. The only hydrodynamic variable is $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989). We consider a nearly flat membrane patch to span at $z=0$ along the XY plane. Assuming simple relaxational dynamics, the general equation of motion of $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989), reads
$\frac{\partial h}{\partial t}=-\frac{\delta{\cal F}}{\delta h}+v_{\text{perm}}+f.$ (3)
Here, ${\cal F}$ is a free energy of an equilibrium fluid membrane, assumed tensionless, and $v_{\text{perm}}$ is the permeation velocity locally normal to the plane of the membrane. Here, ${\cal F}=\int d^{2}x[\frac{K}{2}(\nabla^{2}h)^{2}+C\nabla^{2}h]$, ignoring anharmonic contributions, is a free energy functional, which governs the relaxation of the membrane to equilibrium in the absence of any activity; here, $K>0$ is the bending stiffness and ${\bm{\nabla}}\equiv(\partial/\partial x,\,\partial/\partial y)$ is the 2D gradient operator; $C$ is the spontaneous curvature that can be generically present in an inversion-asymmetric membrane and makes the membrane curve one way or the other depending upon its sign. In the Monge gauge approximation, valid for a planar membrane patch, the spontaneous curvature term being a total derivative term vanishes and makes no contribution to $\cal F$.
Consider a nearly flat patch of an inversion-asymmetric, anisotropic fluid membrane placed in an isotropic, active suspension of actin filaments Marchetti et al. (2013) grafted normally to the membrane locally. Let $\bf p$ be a three-dimensional unit vector ($p^{2}=1$) that gives the local orientation of the actin filaments and $\bf n$ be the local unit normal to the membrane; ${\bf n}=(\hat{z},-\nabla_{i}h)$ to the linear order in height fluctuations, in the present geometry with $i=x,y$. Locally normal grafting of the filaments implies ${\bf p}\cdot{\bf n}=1$. This yields $p_{j}=-\partial_{j}h\,(j=x,y)$ to the linear order in height fluctuations Sarkar and Basu (2013); Banerjee and Basu (2018). Now $v_{\text{perm}}$ Banerjee and Basu (2018), locally normal to an anisotropic membrane Barabási and Stanley (1995), reads
$v_{\text{perm}}=(v_{0}+\xi_{ij}\partial_{i}p_{j}+\tilde{\xi}_{ij}p_{i}p_{j})\sqrt{1+({\bm{\nabla}}h)^{2}}$ (4)
to the lowest order in small fluctuations $p_{i}$; $i=x,y$. In the cellular context, $v_{0}$ sets the scale of the drift speed of the membrane due to actin filament polymerization Sykes et al. (2010).The remaining terms in (4) give the drift due to local orientation fluctuations. Coefficient matrices $\xi_{ij}$ and $\tilde{\xi}_{ij}$ encode the anisotropy, which may originate from the anisotropy in the actomyosin cortex Mayer et al. (2010), or anisotropic inclusions in the membrane Fošnarič et al. (2005). Expanding and substituting for $p_{i}$ and retaining only the lowest order terms in gradients and nonlinearities, we obtain Eq. (11), where $\lambda_{ij}=\tilde{\xi}_{ij}$ is a measure of the average membrane speed due to fluctuations, and $\nu_{ij}=-\xi_{ij}$ is an anisotropic tension. Structures of these anisotropy parameters should depend explicitly on the specific model representation.
Figure 3: Log-log plots of (a) ${\mathcal{W}_{\text{sat}}}$ versus $\ln L$ ($a_{0}=1$) for $\alpha=0.50$, $\beta=2$ (logarithmic rough phase) and $L=100$, $300$, $500$, $800$, $1000$ and $1600$. Red dashed line represents linear fit to the numerical results displayed on a log-log scale with a slope $\tilde{\chi}=0.43\pm 0.019$. (b) ${\mathcal{W}_{\text{sat}}}$ with $L$ for $\alpha=0.50$, $\beta=0.042$ (algebraic rough phase) and $L=100$, $200$, $300$, $400$ and $500$. Red dashed line represents linear fit to the numerical results displayed on a log-log scale with a slope $\chi=0.379\pm 0.008$ (see text).
We now theoretically explain the above numerical observations, quantitatively characterizing the phases, by using Eq. (11). Equation (11) admits a pseudo-Galilean invariance under the transformation $t^{\prime}=t,\,x_{i}^{\prime}=x_{i}+\lambda_{ij}u_{j}t,\,h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+u_{i}x_{i}+\frac{\lambda_{ij}}{2}u_{i}u_{j}t$ sup . In (11), $\nu_{ij}$ plays the role of an (active) anisotropic surface tension. With $\nu>0$, linear stability requires that $\nu\pm\overline{\nu}\geq 0$, or $|\alpha|\leq 1$, i.e. $-1\leq\alpha\leq 1$; otherwise instability ensues. This means $\alpha_{c}=\pm 1$ is a threshold of $\alpha$, with larger $|\alpha|$ implies instability. This is the anisotropy-induced crumpling mentioned earlier in this Letter, which can occur even at zero noises. This instability threshold is independent of wavevector $q$, any ensuing crumpling actually takes place at all scales, meaning an arbitrarily small membrane will be crumpled, provided $|\alpha|>1$. This is reminiscent of the strongly crumpled phase in tensionless equilibrium asymmetric tethered membranes Banerjee et al. (2019a, b). In the crumpled phase, the variance of the local normal fluctuations $\sim\langle({\bm{\nabla}}h)^{2}\rangle$ diverges for a sufficiently large but finite $L$, giving translational and orientational SRO. In this phase, the growth rate is maximum at an angle $\pi/4$ and $43\pi/$ 4, respectively, with the $x$-axis for $\alpha>1$ and $\alpha<-1$. This is reflected in the ripples observed Fig. 2(a) and (b), and the corresponding movies MOV3(a) and MOV3(b) in SM sup .
We now turn to the uncrumpled case with $|\alpha|<1$, and consider its stability and fluctuation properties. In the linear theory with $|\alpha|<1$, width ${\mathcal{W}}_{\text{sat}}\sim\sqrt{\ln(L/a_{0})}$.As in the usual KPZ equation Kardar et al. (1986), the combination of noises and nonlinearities here can change the linear theory results substantially. To study the nonlinear effects in the present model (11), we perform a one-loop perturbative dynamic renormalization group (RG) analysis of Eq. (11) with $|\alpha|<1$. The Wilson dynamic RG method Täuber (2014); Forster et al. (1977) for our model closely resembles that for the KPZ equation Kardar et al. (1986); Barabási and Stanley (1995); Täuber (2014). Galilean invariance of (11) ensures that $\lambda_{ij}$ is unrenormalized, which is an exact statement and the analogue of the nonrenormalization of the nonlinear coupling constant in the usual KPZ equation Barabási and Stanley (1995); Frey and Täuber (1994). However, there are diverging one-loop corrections to $\nu,\,\overline{\nu}$ and $D$, which are represented by the one-loop Feynman diagrams (see SM) sup . Dimensional analysis allows us to identify an effective dimensionless coupling constant $g$ having a critical dimension 2 and a dimensionless ratio $\beta$ defined as
$g=\frac{\lambda^{2}D}{8\pi(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}},\;\beta=\frac{\overline{\lambda}}{\lambda},$ (5)
along with $\alpha$ defined earlier. The RG differential recursion relations ($\exp(l)$ is a dimensionless length) are
$\frac{dD}{dl}=D\bigg{[}z-2-2\chi+g\{\frac{(2+\alpha^{2})+\beta^{2}(1+2\alpha^{2})-6\alpha\beta}{2(1-\alpha^{2})}\}\bigg{]},$ (6)
$\frac{d\nu}{dl}=\nu[z-2-g\{\frac{\alpha\beta}{2}-\frac{\beta^{2}}{2}\}],$ (7)
$\frac{d\overline{\nu}}{dl}=\overline{\nu}[z-2-g\{\frac{1}{2}-\frac{\beta}{2\alpha}\}],$ (8)
$\frac{dg}{dl}=-g^{2}(\beta^{2}-1),$ (9)
together with $d\beta/dl=d/dl(\overline{\lambda}/\lambda)=0$. Here, $z$ and $\chi$ are the dynamic and roughness exponents, respectively Barabási and Stanley (1995)In the linear theory, i.e., with $\lambda_{ij}=0$, $z=2,\,\chi=0$, giving translational QLRO. Equation (9) gives that if $\beta^{2}>1$, then $g(l)\sim\frac{1}{(\beta^{2}-1)l}$ for large “RG time” $l$. On the other hand, if $\beta^{2}<1$, $g(l)$ diverges as $l\rightarrow l_{c}\sim{\cal O}(1)$ from below, where $l_{c}$ is a nonuniversal model parameter-dependent length. Of course, we cannot follow all the way to $g(l)$ diverging; as soon as $g(l)$ becomes ${\cal O}(1)$, the perturbation theory breaks down. This is the strong coupling algebraic rough phase or the model, whose scaling properties are not amenable to perturbative RG treatments, and is analogous to the strong coupling rough phase in the 2D KPZ equation Barabási and Stanley (1995); Tang et al. (1990).
The RG recursion relation of $\alpha$ reads
$\frac{d\alpha}{dl}=\frac{g}{2}\bigr{[}\alpha^{2}\beta-\alpha\beta^{2}-\alpha+\beta\bigr{]}.$
Setting $\frac{d\alpha}{dl}=0$, we obtain
$\alpha^{*}=1/\beta,\beta$
as the fixed points of $\alpha$. Thus in the perturbatively accessible region $\beta^{2}>1$, $\alpha^{*}=1/\beta$, $(\beta)$ is stable (unstable) fixed line. Focusing on the stable case $\beta^{2}>1$ and using the linear theory results $z=2$, $\chi=0$ we find $\nu(l)\sim\nu_{0}\sqrt{l},\,\overline{\nu}(l)\sim\overline{\nu}_{0}\sqrt{l},\,D(l)\sim D_{0}\sqrt{l}$ Mukherjee and Basu (2024), with $\nu_{0},\overline{\nu}_{0},D_{0}$ being the “initial” values of $\nu(l),\,\overline{\nu}(l),D(l)$ respectively. This gives logarithmically faster relaxation than ordinary diffusion, time-scale $t\sim r^{2}/\sqrt{\ln(r/a_{0})}$, where $r$ is a distance, meaning logarithmic breakdown of dynamic scaling. Similar logarithmically modulated noise strength and damping are also found in Refs. Haldar et al. (2023a, b); Jana et al. (2024). We further get
${\mathcal{W}}_{\text{sat}}\sim(D_{0}/{\mathcal{A}})\sqrt{\ln{L}/{a_{0}}},$ (10)
at the stable RG fixed point, where ${\mathcal{A}}\equiv[{4\pi\nu_{0}({1-(\frac{\overline{\nu}_{0}}{\nu_{0}})^{2}})^{1/2}}]^{1/2}$. This corresponds to a logarithmically rough surface with translational QLRO, statistically identical to a 2D Edward-Wilkinson (EW) surface Barabási and Stanley (1995); Edwards and Wilkinson (1982). An effective temperature $T_{\text{eff}}$ defined via $T_{\text{eff}}=D(l)/\nu(l)=D_{0}/\nu_{0}$, i.e., unchanged from its microscopic value. Nonetheless, due to its logarithmically modulated noise and damping, at this RG fixed point our model is different from the 2D EW equation. While the equal-time height-difference correlator is same as in the EW equation, the more general time-displaced correlator is different. With $\alpha^{*}=1/\beta$ and $\beta^{2}>1$ at the RG fixed point, the logarithmic rough phase is macroscopically anisotropic.
In the perturbatively inaccessible case with $\beta^{2}<1$, we have $\alpha^{*}=\beta$ as the stable fixed point, implying large-scale isotropy. In this case $d\nu/dl,d\overline{\nu}/dl=0,dD/dl>0$, reminiscent of the RG flow equations in the 2D isotropic KPZ equation Barabási and Stanley (1995).
Figure 4: (a) Different phases and the RG flow lines of $\alpha$ in $\beta$-$\alpha$ plane. Yellow regions are the logarithmic rough phase. Central white region is the KPZ like putative algebraic rough phase. The black lines in the rough phase show the locus of the fixed point value $\alpha^{*}$ as a function of $\beta$. (b) The RG flow diagram in the $\beta$-$\alpha$-$g$ space. Arrows pointing toward (away from) the $\beta$-$\alpha$ plane represent the direction of RG flows, indicating stable (unstable) regions. Green regions in (a) and (b) correspond to crumpled phase (see text).
Figure 4(a) shows the phases and the RG flow lines of $\alpha$ in $\beta$-$\alpha$ plane. Since $d\beta/dl=0$, the RG flow lines in the $\beta$-$\alpha$ plane are perpendicular to the $\beta$-axis. The green horizontal regions at the top and bottom of Fig. 4(a) have $|\alpha|>1$, corresponding to the crumpled phase. The remaining region of this figure with $|\alpha|<1$, corresponding to linear stability, has two parts: The regions with $\beta^{2}>1$ (marked yellow) are perturbatively accessible logarithmically rough phase, whereas the region with $\beta^{2}<1$ is perturbatively inaccessible. Figure 4(b) shows the RG flow lines in the $\beta$-$\alpha$-$g$ space. Flow lines with initial values of $g$ in the yellow region of $\alpha$-$\beta$ plane flow to zero giving logarithmic roughness. Flow lines with initial values of $g$ inside the whitish cuboid region flow out to the perturbatively inaccessible strong coupling fixed point indicating KPZ-like algebraic roughness.
In our DNS studies, used to corroborate our RG results and characterize the phases, we discretize Eq. (11) by standard forward-backward differences on a square grid with lattice spacing $a_{0}=1$, and numerically solve it by using the Euler algorithm; see SM sup for more details. We calculate ${\mathcal{W}}$ for model parameters falling in the (i) yellow, stable region (logarithmically rough), (ii) unstable, white region (speculated to be algebraically rough) and (iii) green, linearly unstable (crumpled) region in the phase diagram 4(a). For the logarithmic rough phase, we have set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$ and $\overline{\lambda}=4$ such that $|\alpha|<1$ and $|\beta|>1$ and system sizes $L=100$, $300$, $500$, $800$, $1000$ and $1600$ and time step $\Delta t=0.05$ in our simulations. Numerical measurements of ${\cal W}_{\text{sat}}$ for various $L$ give $\tilde{\chi}\approx 0.43\pm 0.019$, close to the analytically obtained value $\tilde{\chi}=0.50$, firmly establishing the logarithmic rough phase.
Next we set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$ and $\overline{\lambda}=1$ such that $|\alpha|<1$ and $|\beta|<1$, corresponding to $dg(l)/dl>0$ [see Eq. (9)]. Time step $\Delta t=0.01$. With ${\cal W}\sim t^{\tilde{\beta}}$ initially, we obtained $\tilde{\beta}\approx 0.234\pm 1.644\times 10^{-5}$; see also SM sup . Furthermore, numerically measuring ${\mathcal{W}}_{\text{sat}}$ for $L=100$, $200$, $300$, $400$ and $500$ give $\chi\approx 0.379\pm 0.008$, and hence $z=\chi/\tilde{\beta}=1.619$ with $\chi+z\approx 2$ (as it should be). Our results on $\chi$ confirms an algebraically rough phase with translational SRO and orientational LRO, as speculated from our RG calculations above. The value of $\chi=0.379$ shows reasonable agreement with the existing results Oliveira et al. (2013); Aarão Reis (2004); Kondev et al. (2000); Pagnani and Parisi (2015); Kelling et al. (2017); Ódor et al. (2009); Halpin-Healy (2012); Chin and den Nijs (1999); Colaiori and Moore (2001); Gomes-Filho et al. (2021) and also with a mode coupling theory prediction Bhattacharjee (1998), indicating scaling behavior statistically same as in the strong coupling rough phase of the 2D KPZ equation. These DNS results lend further credence to a strong coupling rough phase with large-scale isotropy.
So far we have considered a nearly flat patch within a Monge parametrization, which holds up to length scales at which curvature $C$ becomes important. Within this scale, while microscopically the membrane is inversion-asymmetric, the effective long wavelength behavior is symmetric in the logarithmic rough phase. The nonlinear terms with coefficients $\lambda_{ij}$ break the inversion symmetry. However, coupling $g$ vanishes in the long wavelength limit (albeit slowly enough to introduce logarithmic modulations to the noise and damping). At larger scales, when $C$ should be important, the spontaneous curvature term scales as $C/r$, whereas the surface tension term scales as $r^{0}$, making $C$ increasingly unimportant (relative to the surface tension) at larger length scales. We thus expect the membrane to remain effectively inversion-symmetric (and statistically flat) at arbitrarily large scales. In the algebraic rough phase, $g$ is not expected to vanish, giving inversion-asymmetry to survive in the long wavelength limit. Nonetheless, with expected $z<2$ the surface tension continues to dominate over $C$, ultimately giving a statistically flat but inversion-asymmetric membrane. At the onset of the crumpled phase, the effective surface tension vanishes, making $C$ important. This is beyond the scope of the present study.
We have thus shown that the hydrodynamic theory for a moving active anisotropic inversion-asymmetric fluid membrane is given by an anisotropic KPZ equation, giving a logarithmically rough phase an algebraically rough phase and a crumpled phase. The average membrane velocity $\lambda_{ij}\langle{\nabla_{i}h}\nabla_{j}h\rangle$, an observable feature and proportional to $g$, vanishes in the inversion-symmetric logarithmic rough phase, but is finite in the asymmetric algebraically rough phase. It would be interesting to explore connections between our predicted crumpled phase and observed reduction of membrane tension to even negative values during membrane fusion Solon et al. (2006). Membrane tension may be measured using optical set ups Popescu et al. (2006), giving information about $\nu,\overline{\nu}$. Comparison with experimental results on membrane fluctuations by using, e.g., spectroscopy Betz and Sykes (2012) can yield information about $\lambda,\overline{\lambda}$. Studying how the fluid flows along the membrane in the different phases and their dependence on the tension anisotropy are related to the cortical flows reported in Ref. Mayer et al. (2010) should be promising future work directions.
For reasons of simplicity, we set $\nu_{xx}=\nu_{yy}$ and $\lambda_{xx}=\lambda_{yy}$. We note that $\nu_{xx}\neq\nu_{yy}$ together with $\lambda_{xx}\neq\lambda_{yy}$ (but without $\overline{\nu},\overline{\lambda}$) was considered in Refs. Wolf (1991); Altman et al. (2015), giving QLRO translational order and 2D KPZ-like instability for different choices of an effective anisotropy parameter. We thus speculate that even in the general case with $\nu_{xx}\neq\nu_{yy}$, $\lambda_{xx}\neq\lambda_{yy}$ and with non-vanishing $\overline{\nu},\overline{\lambda}$ can still have just the three phases studied here.It will be interesting to connect Eq. (11) to variants of the Toom model Barabási et al. (1992); Jeong et al. (1993), and generalize our study by considering the interplay between the nonlocal effects Jana et al. (2024); Jana and Basu (2024) and anisotropy.
Acknowledgment:- A.H. thanks Alexander von Humboldt (AvH) Stiftung (Germany) for a postdoctoral fellowship. A.B. thanks the AvH Stiftung (Germany) for partial financial support under the Research Group Linkage Programme scheme (2024).
Appendix A Linear instability analysis
We start from Eq. (1) of the main text for an anisotropic membrane with a fixed background, ignoring the conservation of embedding bulk fluid momentum. As argued, the hydrodynamic equation governing the undulation modes of such a membrane, described by the height field $h({\bf x},t)$, is an anisotropic Kardar-Parisi-Zhang (KPZ) equation with a Gaussian-distributed, zero-mean white noise $f({\bf x},t)$
$\frac{\partial h}{\partial t}+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (11)
Here, both $i$ and $j$ can take value $1$ or $2$. To perform linear stability analysis, we consider only the linear terms in Eq. (11), i.e., set $\lambda_{ij}=0$. The resulting linearized equation of $h$ in Fourier space is
$\partial_{t}h=-\nu(k_{x}^{2}+k_{y}^{2}+2\alpha k_{x}k_{y})h,$ (12)
[…]
When $(1+\alpha\sin 2\theta)<0$, linear instability ensues. The growth rate of initial fluctuations is maximized when $(1+\alpha\sin 2\theta)$ attains its largest negative value. For $\alpha>1$, this happens at $\theta=3\pi/4$, corresponding to $k_{x}=-k_{y}$ giving the Fourier wavevector with the maximum growth rate. Keeping only the corresponding Fourier mode (which is the most dominant Fourier mode), we get
$h(\mathbf{r})\approx h_{\mathbf{k}}e^{i\frac{k}{\sqrt{2}}(-x+y)},$ (13)
which has a phase given by
$\phi(x,y)=\frac{k}{\sqrt{2}}(y-x).$ (14)
If we move in the $xy$-plane along paths with a constant phase, i.e., with constant $y-x$, the exponential factor $e^{i\phi}$ does not change. Consequently, the height function $h(x,y)$ does not vary along these directions. These lines, defined by $y-x=\text{constant}$, correspond to the stripes observed in real space, forming at an angle of $\pi/4$. The periodicity of the stripes is controlled by the wavenumber $k_{\text{max}}\sim\pi/a_{0}\sim\pi$, given that $a_{0}=1$ in our simulations. A representative snapshot of this pattern can be seen in Fig. 2(a) of the main text. On the other hand, for $\alpha<-1$, the maximum growth rate occurs at $\theta=\pi/4$, corresponding to $k_{x}=k_{y}$. Now using the logic used above, we find this corresponds to the formation of stripes at an angle of $43\pi/$ 4 in the $x$-$y$ plane. A representative snapshot of this case is shown in Fig. 2(b) of the main text.
Appendix B Galilean invariance
We now explicitly show that under the Galilean transformation defined by
$t\rightarrow t^{\prime}\equiv t,\,x_{i}\rightarrow x_{i}^{\prime}\equiv x_{i}+\lambda_{ij}u_{j}t,\,h({\bf x},t)\rightarrow h^{\prime}({\bf x}^{\prime},t^{\prime})\equiv h({\bf x},t)+u_{i}x_{i}+\frac{\lambda_{ij}}{2}u_{i}u_{j}t,$ (15)
Eq. (11) remains invariant. Under the Galilean transformation
$\frac{\partial}{\partial t}\rightarrow\frac{\partial}{\partial t^{\prime}}+\lambda_{ij}u_{j}\frac{\partial}{\partial x_{i}^{\prime}},\,\frac{\partial}{\partial x_{i}}\rightarrow\frac{\partial}{\partial x_{i}^{\prime}},$
$\frac{\partial h({\bf x},t)}{\partial t}\rightarrow\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial t^{\prime}}+\lambda_{ij}u_{j}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x_{i}^{\prime}}-\frac{\lambda_{ij}}{2}u_{i}u_{j}.$ (16)
$\frac{\lambda_{ij}}{2}\frac{\partial h({\bf x},t)}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{j}}\rightarrow\frac{\lambda_{ij}}{2}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{j}}+\frac{\lambda_{ij}}{2}u_{i}u_{j}-\lambda_{ij}u_{j}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}.$ (17)
$\nu_{ij}\frac{\partial}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{j}}\rightarrow\nu_{ij}\frac{\partial}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{j}}.$ (18)
The variance of the noise $f$, being $\delta$-correlated in time, remains invariant under this transformation. Using Eq. (16), (17), (18) and rewriting Eq. (11) in terms of $h^{\prime}({\bf x^{\prime}},t^{\prime})$, ${\bf x^{\prime}}$ and $t^{\prime}$, we find that Eq. (11) is Galilean invariant.
Appendix C Renormalization group calculations
We start by constructing a generating functional Bausch et al. (1976); Täuber (2014) by using Eq. (1) and Eq. (2) of the main text. We get
$\mathcal{Z}=\int\mathcal{D}\hat{h}\mathcal{D}he^{-\mathcal{S}[\hat{h},h]},$ (19)
where $\hat{h}$ is referred as response field and $\mathcal{S}$ is action functional, which is given by
$S=-\int_{x,t}\hat{h}D\hat{h}+\int_{x,t}\hat{h}(\partial_{t}h+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h-\nu_{ij}\partial_{i}\partial_{j}h).$ (20)
C.1 Results from the linearized equation of motion
This can be obtained from the Gaussian part of the action functional (20), i.e., by setting $\lambda_{ij}=0$ in (20).
We first define Fourier transforms in space and time by
$h({\bf x},t)=\int_{{\bf k},\omega}h({\bf k},\omega)e^{i({\bf k\cdot x}-\omega t)}.$
Then the (bare) two-point functions in the Gaussian limit of (20) are
$\langle\hat{h}({\bf k},\omega)\hat{h}(-{\bf k},-\omega)\rangle_{0}=0,$ (21a)
$\langle\hat{h}({\bf k},\omega)h(-{\bf k},-\omega)\rangle_{0}=\frac{1}{i\omega+\nu_{ij}k_{i}k_{j}},$ (21b)
$\langle\hat{h}(-{\bf k},-\omega)h({\bf k},\omega)\rangle_{0}=\frac{1}{-i\omega+\nu_{ij}k_{i}k_{j}},$ (21c)
$\langle h({\bf k},\omega)h(-{\bf k},-\omega)\rangle_{0}=\frac{2D}{\omega^{2}+(\nu_{ij}k_{i}k_{j})^{2}}.$ (21d)
Figure 5: Diagrammatic representations of two point functions.
Figure 5 shows the diagrammatic representations of the various two-point functions.
C.2 One-loop corrections to the model parameters
In the (more usual) isotropic models, applications of the Wilson momentum space RG involves integrating the wavevector magnitude $q$ within a shell $\Lambda/b\leq q\leq\Lambda,\,b>1$ Hohenberg and Halperin (1977); Täuber (2014). The present model is however anisotropic, lacking any spherical symmetry. Thus, using spherical coordinates and a spherical shell for mode elimination is not convenient for the present problem. Instead, we integrate $q_{x}$ over the whole range from $\infty$- ∞ to $\infty$, and then integrate $q_{y}$ over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$ with $b>1$; see Fig. 6. Of course, we could have chosen another integration scheme, in which $q_{y}$ is integrated over the whole range from $\infty$- ∞ to $\infty$, and then integrate $q_{x}$ over two thin strips $\Lambda/b\leq q_{x}\leq\Lambda$ and $-\Lambda\leq q_{x}\leq-\Lambda/b$. For our model, the two schemes are equivalent, connected to the fact that the different anisotropic terms of the same order scale in the same way.
[…]
Figure 6: Wavevector-shell integration in the one-loop diagrams obtained by perturbative expansions of the action functional (20), corresponding to the anisotropic KPZ equation. Here, $q_{x}$ is integrated over the range $\infty$- ∞ to $\infty$, while $q_{y}$ is integrated over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$.
We now proceed with the evaluation of the integrals that arise from one-loop Feynman diagrams. These integrals take the form
$\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{m}}=\frac{1}{\nu^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}^{2}+2\alpha q_{x}q_{y}+q_{y}^{2})^{m}}$
$=\frac{1}{\nu^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}-a_{+})^{m}(q_{x}-a_{-})^{m}},$ (22)
where $\tilde{f}(q_{x},q_{y})$ represents functions of $q_{x}$ and $q_{y}$, and we use $\alpha={\overline{\nu}}/{\nu}$. The exponent $m$ takes values $2$ or $3$. The terms $a_{+}=q_{y}(-\alpha+ip)$ and $a_{-}=q_{y}(-\alpha-ip)$ are two poles of order $m$, where $p=\sqrt{1-\alpha^{2}}$.
To evaluate this integral, we first perform the integration over $q_{x}$ in the range (,)($\infty$,$\infty$)( - ∞ , ∞ ). Subsequently, we integrate over the $y$-component of momentum within the two high-$|q_{y}|$ strips, specifically in the regions $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$. See Fig. 6. For computational convenience, we define $q_{x}^{\prime}=q_{x}+\alpha q_{y}$. Under this change of variables, Eq. (22) transforms into
$\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}+\int_{-\Lambda}^{-\Lambda/b}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}.$ (23)
The first integral receives contributions from the positive values of $q_{y}$, while the second integral accounts for the negative values of $q_{y}$. Since the denominator of each of the integrals is even in $q_{x}^{\prime}$, any terms in $\tilde{f}(q_{x}^{\prime},q_{y})$ containing odd powers of $q_{x}^{\prime}$ must vanish under integration over $q_{x}^{\prime}$. Only terms with even powers of $q_{x}^{\prime}$ survive and contribute to the integral. It turns out from the explicit expressions of the various one-loop integrals that these are also even powers of $q_{y}$. Consequently, both integrals in Eq. (23) yield the same result. Hence, in the following analysis, we evaluate the integrals considering only the positive values of $q_{y}$, noting that the contributions from negative $q_{y}$ are identical.
Below we now evaluate the one-loop diagrams that account for fluctuation-induced corrections to the model parameters.
C.3 One-loop corrections to $D$
The one-loop diagram correcting $D$ is given in Fig. 7, which comes with a symmetry factor of $2$.
Figure 7: One-loop Feynman diagram that contributes to the correction of $D$.
It is given by
$\lambda_{ij}\lambda_{mn}\int\frac{d\Omega}{2\pi}\frac{d^{2}{q}}{(2\pi)^{2}}(q_{i})(k-q)_{j}\,q_{m}\,(k-q)_{n}\frac{D^{2}}{\Big{[}\Omega^{2}+(\nu_{lk}q_{l}q_{k})^{2}\Big{]}\Big{[}(\omega-\Omega)^{2}+\big{(}\nu_{ps}(k-q)_{p}(k-q)_{s}\big{)}^{2}\Big{]}}.$ (24)
Setting $k=0$, $\omega=0$, the above integral reduces to
$\lambda_{ij}\lambda_{mn}\int_{q,\Omega}\frac{q_{i}q_{j}q_{m}q_{n}D^{2}}{\Big{[}\Omega^{2}+(\nu_{xx}q_{x}^{2}+2\nu_{xy}q_{x}q_{y}+\nu_{yy}q_{y}^{2})^{2}\Big{]}^{2}}.$ (25)
After performing the $\Omega$-integral, the contribution in (25) becomes
$\frac{\lambda_{ij}\lambda_{mn}}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{D^{2}q_{i}q_{j}q_{m}q_{n}}{(\nu_{xx}q_{x}^{2}+2\nu_{xy}q_{x}q_{y}+\nu_{yy}q_{y}^{2})^{3}}.$ (26)
Using $\lambda_{xx}=\lambda_{yy}=\lambda$ and $\lambda_{xy}=\lambda_{yx}=\overline{\lambda}$ we get
$\lambda_{ij}\lambda_{mn}q_{i}q_{j}q_{m}q_{n}=\lambda^{2}q_{x}^{4}+\lambda^{2}q_{y}^{4}+4\overline{\lambda}^{2}q_{x}^{2}q_{y}^{2}+4\lambda\overline{\lambda}q_{x}^{3}q_{y}+4\lambda\overline{\lambda}q_{y}^{3}q_{x}+2\lambda^{2}q_{x}^{2}q_{y}^{2}.$
Substituting this in Eq. (26) we get the correction to $D$ as
$\frac{D^{2}}{4\nu^{3}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}}{2\pi}\Biggl{[}\frac{\lambda^{2}q_{x}^{4}+(2\lambda^{2}+4\overline{\lambda}^{2})q_{x}^{2}q_{y}^{2}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}+\frac{4\lambda\overline{\lambda}q_{x}^{3}q_{y}+4\lambda\overline{\lambda}q_{x}q_{y}^{3}+\lambda^{2}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}\Biggl{]},$ (27)
where $a_{+}=q_{y}[-\alpha+ip]$ and $a_{-}=q_{y}[-\alpha-ip]$ are two poles of third order. We set $\nu_{xx}=\nu_{yy}=\nu$, $\nu_{xy}=\nu_{yx}=\overline{\nu}$, and $p=\sqrt{1-\alpha^{2}}$. Also defining
$\mathcal{A}=\lambda^{2},\mathcal{B}=(2\lambda^{2}+4\overline{\lambda}^{2}),\mathcal{C}=4\lambda\overline{\lambda},\mathcal{D}=4\lambda\overline{\lambda},\mathcal{E}=\lambda^{2},$
Equation (27) becomes
$\frac{D^{2}}{4\nu^{3}}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}.$ (28)
Now we perform the $q_{x}$-integral first. Considering a closed contour in the upper half complex plane, the relevant poles within the closed contour are $q_{x}=a_{+}=q_{y}(-\alpha+ip)$. The residue at $q_{x}=a_{+}$ is
$\frac{1}{2!}\frac{d^{2}}{dq_{x}^{2}}\biggr{[}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{-})^{3}}\biggr{]}_{q_{x}=a_{+}}.$
Evaluating, this gives
$\frac{1}{2}\biggr{[}\frac{12\mathcal{A}q_{x}^{2}+2\mathcal{B}q_{y}^{2}+6\mathcal{C}q_{x}q_{y}}{(q_{x}-a_{-})^{3}}-6\frac{4\mathcal{A}q_{x}^{3}+2\mathcal{B}q_{x}q_{y}^{2}+3\mathcal{C}q_{x}^{2}q_{y}+\mathcal{D}q_{y}^{3}}{(q_{x}-a_{-})^{4}}+12\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{-})^{5}}\biggr{]}_{q_{x}=a_{+}}.$
Substituting $q_{x}=a_{+}$ in above expression and multiplying with $2\pi i$ gives the result for the $q_{x}$-integral. Thus Eq. (28) becomes
$\frac{3D^{2}}{64\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\mathcal{A}+\mathcal{E})+\mathcal{B}\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-(\mathcal{C}+\mathcal{D})\alpha\biggr{]}.$ (29)
After performing the $q_{y}$-integral and using the values of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ Eq. (29) can be written as
$\frac{3D^{2}}{16\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\frac{\ln b}{2\pi}\biggr{[}\frac{\lambda^{2}}{4}\Bigl{(}2+\frac{2}{3}+\frac{4}{3}\alpha^{2}\Bigl{)}-2\lambda\overline{\lambda}\alpha+\overline{\lambda}^{2}\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}\biggr{]}.$ (30)
Using $\alpha={\overline{\nu}}/{\nu}$, Eq. (30) becomes
$\frac{3D^{2}}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\biggr{[}\frac{\lambda^{2}}{4}\Bigl{(}\frac{8\nu^{2}}{3(\nu^{2}-\overline{\nu}^{2})}+\frac{4\overline{\nu}^{2}}{3(\nu^{2}-\overline{\nu}^{2})}\Bigl{)}+\overline{\lambda}^{2}\Bigl{(}\frac{\nu^{2}}{3(\nu^{2}-\overline{\nu}^{2})}+\frac{2\overline{\nu}^{2}}{3(\nu^{2}-\overline{\nu}^{2})}\Bigl{)}-\frac{2\lambda\overline{\lambda}\nu\overline{\nu}}{(\nu^{2}-\overline{\nu}^{2})}\biggr{]}.$ (31)
C.4 One-loop corrections to $\nu_{ij}$
The one-loop Feynman diagram, with a symmetry factor 8, that contributes to the fluctuation correction of $\nu$ and $\tilde{\nu}$ is shown in Fig. 8.
Figure 8: One-loop Feynman diagram that contributes to the correction of $\nu$ and $\overline{\nu}$.
It is given by
$-\lambda_{ij}\lambda_{mn}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}q_{m}k_{n}\langle|h_{{\bf q},\Omega}|^{2}\rangle\langle h_{{\bf k-q},-\Omega}\hat{h}_{{\bf-k+q},\Omega}\rangle.$
After symmetrizing, i.e., ${\bf q}\rightarrow{{\bf q}+{{\bf k}}/{2}}$, and performing the $\Omega$-integral, we are left with two integrals that include corrections for $\nu$ and $\overline{\nu}$. These integrals are
$\mathcal{I}_{1}=\frac{\lambda_{ij}\lambda_{mn}k_{n}k_{m}D}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q_{i}q_{j}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}},$ (32)
and
$\mathcal{I}_{2}=-\frac{\lambda_{ij}\lambda_{mn}k_{n}D}{2}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q_{i}q_{j}q_{m}\bigl{[}\nu(k_{x}q_{x}+k_{y}q_{y})+\overline{\nu}(k_{x}q_{y}+k_{y}q_{x})\bigl{]}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{3}}.$ (33)
Corrections to $\nu$ are given by the coefficient of $k_{x}^{2}$ (or $k_{y}^{2}$). Considering the coefficient of $k_{x}^{2}$, $\lambda_{ij}\lambda_{mn}k_{n}k_{m}q_{i}q_{j}=k_{x}^{2}\lambda(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})$. Substituting this in Eq. (32), the correction to $\nu$ coming from $\mathcal{I}_{1}$ is
$\frac{D\lambda k_{x}^{2}}{4}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}}=\frac{D\lambda k_{x}^{2}}{4\nu^{2}}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{(q_{x}-a_{+})^{2}(q_{x}-a_{-})^{2}}.$ (34)
To evaluate the $q_{x}$-integral, considering a closed contour in the upper half plane the relevant pole $q_{x}=a_{+}=q_{y}(-\alpha+ip)$, which is a second order pole. The residue at $q_{x}=a_{+}$ is
$\frac{d}{dq_{x}}\Biggl{[}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{(q_{x}-a_{-})^{2}}\Biggl{]}_{q_{x}=a_{+}}=-i\frac{\lambda-\overline{\lambda}\alpha}{2p^{3}q_{y}}.$
Multiplying by $2\pi i$ and writing $p$ in terms of $\alpha$, Eq. (34) becomes
$k_{x}^{2}\frac{\lambda D(\lambda-\overline{\lambda}\alpha)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{(2\pi)q_{y}}.$ (35)
Equation (35) gives the correction to $\nu$ coming from $\mathcal{I}_{1}$. Similarly keeping only order of $k_{x}^{2}$, Eq. (33) becomes
$-k_{x}^{2}\frac{D}{2\nu^{3}}\int_{q_{x},q_{y}}\biggr{[}\frac{\lambda^{2}\nu q_{x}^{4}+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})q_{x}^{2}q_{y}^{2}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}+\frac{(3\lambda\overline{\lambda}\nu+\lambda^{2}\overline{\nu})q_{x}^{3}q_{y}+(\lambda^{2}\overline{\nu}+2\overline{\lambda}^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)q_{x}q_{y}^{3}+\lambda\overline{\lambda}\overline{\nu}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}\biggr{]}.$ (36)
This integral is similar to the integral in Eq. (27). However, $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ are now different with values:
$\mathcal{A}=\lambda^{2}\nu,\mathcal{B}=(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu}),\mathcal{C}=3\lambda\overline{\lambda}\nu+\lambda^{2}\overline{\nu},\mathcal{D}=\lambda^{2}\overline{\nu}+2\overline{\lambda}^{2}\overline{\nu}+\lambda\overline{\lambda}\nu,\mathcal{E}=\lambda\overline{\lambda}\overline{\nu},$
Comparing with Eq. (29) and substituting these values of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ equation (36) can be written as
$-k_{x}^{2}\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\Biggl{[}(\lambda^{2}\nu+\lambda\overline{\lambda}\overline{\nu})-(2\lambda^{2}\overline{\nu}+4\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu})\alpha+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})\Bigr{(}\frac{1+2\alpha^{2}}{3}\Bigr{)}\Biggl{]}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (37)
Equation (37) gives the correction to $\nu$ coming from $\mathcal{I}_{2}$. Adding Eqs. (35) and (37) and extracting the coefficient of $k_{x}^{2}$, we get the total correction to $\nu$ at the one-loop order as
$-\frac{3D}{32\nu^{3}{(1-\alpha^{2})^{\frac{5}{2}}}}\frac{\ln b}{2\pi}\Biggl{[}(\lambda^{2}\nu+\lambda\overline{\lambda}\overline{\nu})-(2\lambda^{2}\overline{\nu}+4\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu})\alpha+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})\Bigr{(}\frac{1+2\alpha^{2}}{3}\Bigr{)}\Biggl{]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{D}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\Biggl{[}\lambda^{2}-\lambda\overline{\lambda}\alpha\Biggl{]}.$ (38)
Substituting $\alpha={\overline{\nu}}/{\nu}$ in above equation and after some algebra the above expression reduces to
$\frac{\nu D}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{5}{2}}}\frac{\ln b}{2\pi}\Biggl{[}(\nu^{2}-\overline{\nu}^{2})\Big{(}\frac{\lambda\overline{\lambda}\overline{\nu}}{\nu}-\overline{\lambda}^{2}\Big{)}\Biggl{]}.$ (39)
To get the corrections to $\overline{\nu}$, we need to find coefficient of $2k_{x}k_{y}$ from Eqs. (32) and (33). Therefore coonsidering only the $k_{x}k_{y}$-term, $\lambda_{ij}\lambda_{mn}k_{n}k_{m}q_{i}q_{j}\rightarrow 2k_{x}k_{y}\overline{\lambda}\bigl{[}\lambda(q_{x}^{2}+q_{y}^{2})+2\overline{\lambda}q_{x}q_{y}\bigl{]}$. Substituting this in Eq. (32) we get
$\frac{D\overline{\lambda}(2k_{x}k_{y})}{4}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}}.$ (40)
The integral in Eq. (40) is same as the integral in Eq. (34). Comparing with Eq. (35) we can write Eq. (40) as
$(2k_{x}k_{y})\frac{D\overline{\lambda}(\lambda-\overline{\lambda}a)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (41)
Equation (41) gives the correction to $\overline{\nu}$ that comes from $\mathcal{I}_{1}$. Now,
$\lambda_{ij}\lambda_{mn}k_{n}q_{i}q_{j}q_{m}=k_{x}(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})(\lambda q_{x}+\overline{\lambda}q_{y})+k_{y}(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})(\lambda q_{y}+\overline{\lambda}q_{x}).$
Substituting this in Eq. (33) and extracting the coefficient of $2k_{x}k_{y}$ we get
$-(2k_{x}k_{y})\frac{D}{2\nu^{3}}\int_{q_{x},q_{y}}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}.$ (42)
Notice that the integral in Eq. (42) has the same form as the integral in Eq. (28), but now $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ are given by $\mathcal{A}=\mathcal{E}=\bigr{(}\frac{\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu}{2}\bigr{)},\mathcal{B}=(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu}),\mathcal{C}=\mathcal{D}=(\lambda^{2}\nu+\overline{\lambda}^{2}\nu+2\lambda\overline{\lambda}\overline{\nu})$. Thus comparing with Eq. (29) we write Eq. (42) as
$-(2k_{x}k_{y})\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)+(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu})\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-2(\lambda^{2}\nu+\overline{\lambda}^{2}\nu+2\lambda\overline{\lambda}\overline{\nu})\alpha\biggr{]}.$ (43)
Equation (43) gives the correction to $\overline{\nu}$ that comes from $\mathcal{I}_{2}$. Thus adding Eq. (41) and (43) and extracting the coefficient of $2k_{x}k_{y}$, we get total correction to $\overline{\nu}$ given by
$-\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)+(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu})\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-(2\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+4\lambda\overline{\lambda}\overline{\nu})\alpha\biggr{]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{D(\overline{\lambda}\lambda-\overline{\lambda}^{2}\alpha)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (44)
Substituting $\alpha={\overline{\nu}}/{\nu}$ in the above expression and after some algebraic simplification the above contribution becomes
$\frac{\lambda^{2}\overline{\nu}D}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\Biggl{[}1-\frac{\overline{\lambda}\nu}{\lambda\overline{\nu}}\Biggl{]}.$ (45)
C.5 Nonrenormalization of $\lambda$ and $\overline{\lambda}$
Figure 9: One-loop Feynman diagrams giving the corrections to $\lambda$ and $\overline{\lambda}$.
The left diagram in Fig. 9 has a symmetry factor 24 and includes corrections to both $\lambda$ and $\overline{\lambda}$. We find
$-\lambda_{ij}\lambda_{mn}\lambda_{rs}D\int_{q,\Omega}\frac{(q_{i}q_{j})(-k_{1m}q_{n})(q_{r}k_{2s})}{(-i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})(i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})\Bigl{(}\Omega^{2}+(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{)}}$
$=\frac{\lambda_{ij}\lambda_{mn}\lambda_{rs}D}{4}\int_{q}\frac{(q_{i}q_{j})(k_{1m}q_{n})(q_{r}k_{2s})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{3}}.$ (46)
Similarly the right diagram in Fig. 9 has a symmetry factor $48$ and contributes
$-\lambda_{ij}\lambda_{mn}\lambda_{rs}(2D)\int_{q,\Omega}\frac{(q_{i}q_{j})(-k_{1m}q_{n})(-q_{r}k_{2s})}{(-i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{(}\Omega^{2}+(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{)}}$
$=-\frac{\lambda_{ij}\lambda_{mn}\lambda_{rs}D}{4}\int_{q}\frac{(q_{i}q_{j})(k_{1m}q_{n})(q_{r}k_{2s})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{3}}.$ (47)
Equations (46) and (47), when added, give zero contribution to the correction of $\lambda$ and $\overline{\lambda}$. These are consistent with the Galilean invariance of Eq. (1) in the main text.
C.6 RG recursion relations
We integrate out the dynamical fields with wavevectors $k_{y}$ in the range ($\Lambda/b\leq k_{y}\leq\Lambda$ and $-\Lambda\leq k_{y}\leq-\Lambda/b$) in the action functional (20). Now we rescale space, time and fields to raise the wavevector cutoff of $k_{y}$ to ±$\pm\Lambda$. We rescale ${\bf k}\to{\bf k}^{{}^{\prime}}=b{\bf k}\implies{\bf x}\to{\bf x}^{{}^{\prime}}=\frac{{\bf x}}{b}$ similarly, $\omega\to\omega^{{}^{\prime}}=b^{z}\omega\implies t{{}^{\prime}}=\frac{t}{b^{z}}$. Furthermore, the fields are rescaled as
$h^{<}({\bf k},\omega)=\xi h({\bf k}^{{}^{\prime}},\omega^{{}^{\prime}}),$
$\hat{h}^{<}({\bf k},\omega)=\hat{\xi}\hat{h}({\bf k}^{{}^{\prime}},\omega^{{}^{\prime}}).$
Using these and imposing the condition that coefficient of $\int_{{\bf x},t}\hat{h}\partial_{t}h$ remains unity under rescaling we find that, model parameters scale as
$\nu_{ij}^{{}^{\prime}}=\nu_{ij}^{<}b^{z-2},$
$D^{{}^{\prime}}=D^{<}b^{d+3z}\xi^{-2},$
$\lambda_{ij}^{{}^{\prime}}=\lambda_{ij}^{<}b^{-d-2}\xi.$
Also, $h^{<}({\bf x},t)$ scale as $b^{-d-z}\xi h({\bf x}^{{}^{\prime}},t^{{}^{\prime}})$. Defining $\xi_{R}=b^{-d-z}\xi=b^{\chi}$ we get model parameters rescale as
$D^{{}^{\prime}}=D^{<}b^{z-d-2\chi},$
$\nu^{{}^{\prime}}=\nu^{<}b^{z-2},$
$\overline{\nu}^{{}^{\prime}}=\overline{\nu}^{<}b^{z-2},$
$\lambda^{{}^{\prime}}=\lambda^{<}b^{\chi+z-2},$
$\overline{\lambda}^{{}^{\prime}}=\overline{\lambda}^{<}b^{\chi+z-2}.$
Here, $\chi$ is the roughness exponent of an anisotropic KPZ surface. Since there is no correction to $\lambda$, $\overline{\lambda}$ so $\lambda^{<}=\lambda$ and $\overline{\lambda}^{<}=\overline{\lambda}$. We further get
$D^{<}$ $=$ $D\biggr{[}1+\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{[}\Bigl{(}\frac{2\nu^{2}+\overline{\nu}^{2}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}+\frac{\overline{\lambda}^{2}}{\lambda^{2}}\Bigl{(}\frac{\nu^{2}+2\overline{\nu}^{2}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}-\frac{6\overline{\lambda}}{\lambda}\Bigl{(}\frac{\nu\overline{\nu}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}\Bigr{]}\ln b\biggr{]},$ (48)
$\nu^{<}$ $=$ $\nu\biggr{[}1-\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{(}\frac{\overline{\lambda}\overline{\nu}}{\lambda\nu}-(\frac{\overline{\lambda}}{\lambda})^{2}\Bigr{)}\ln b\biggr{]},$ (49)
$\overline{\nu}^{<}$ $=$ $\overline{\nu}\biggr{[}1-\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{(}1-\frac{\overline{\lambda}\nu}{\lambda\overline{\nu}}\Bigr{)}\ln b\biggr{]}.$ (50)
Now set $b=e^{\delta l}\approx 1+\delta l$ for small $\delta l$, and define a dimensionless coupling constant by $g=\frac{\lambda^{2}D}{8\pi(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}$ and dimensionless ratios $\alpha=\frac{\overline{\nu}}{\nu}$ and $\beta=\frac{\overline{\lambda}}{\lambda}$ (see main text). These allow us to obtain the differential RG recursion relations reported in the main text. Using the flow equations of $g$ and $\alpha$ it turns out that $g$ is stable if $\beta^{2}>1$ and in this range of $\beta$, $\alpha^{*}=1/\beta$ is locus of stable fixed points (stable fixed line). Using $d=2$ and the linear theory results $z=2$, $\chi=0$ we find that $\nu_{l}\sim\nu_{0}l^{A}$, $\overline{\nu}_{l}\sim\overline{\nu}_{0}l^{B}$ and $D_{l}\sim D_{0}l^{C}$. Where,
$A[\beta,\alpha^{*}]=\frac{\beta^{2}-\alpha^{*}\beta}{2s},\;B[\beta,\alpha^{*}]=\frac{(\beta/\alpha^{*})-1}{2s},\;C[\beta,\alpha^{*}]=\frac{1}{2s}\biggr{[}\frac{(2+\alpha^{*2})+\beta^{2}(1+2\alpha^{*2})-6\beta\alpha^{*}}{1-\alpha^{*2}}\biggr{]}.$
Here, $s=\beta^{2}-1$. It turns out that at $\alpha=\alpha^{*}=\frac{1}{\beta}$, $A=B=C=\frac{1}{2s}(\beta^{2}-1)=\frac{1}{2}$ implying $\nu_{l}\sim\nu_{0}\sqrt{l}$, $\overline{\nu}_{l}\sim\overline{\nu}_{0}\sqrt{l}$ and $D_{l}\sim D_{0}\sqrt{l}$.
Appendix D Direct numerical simulations
Following Ref. Moser et al. (1991), we discretize Eq. (11) by standard forward-backward differences on a square grid with lattice constant $a_{0}$. Employing the Euler algorithm with time increments $\Delta t$, we integrate Eq. (11). Grid points are labeled by ${\bf n}$, and the basis vector is denoted by ${\bf e}_{i}$, where $i=1,2$, giving the discretized form of Eq. (11)
$h_{\bf n}(t+\Delta t)$ $=h_{\bf n}(t)+\frac{\Delta t}{(a_{0})^{2}}\sum_{i=1}^{2}\biggl{\{}\nu\Bigl{[}h_{\bf n+e_{i}}(t)-2h_{\bf n}(t)+h_{\bf n-e_{i}}(t)\Bigl{]}-\frac{\lambda}{8}\Bigl{[}h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl{]}^{2}\biggl{\}}+\frac{2\overline{\nu}\Delta t}{4(a_{0})^{2}}\Bigl{[}h_{\bf n+e_{1}+e_{2}}(t)$
$-h_{\bf n+e_{1}-e_{2}}(t)+h_{\bf n-e_{1}-e_{2}}(t)-h_{\bf n-e_{1}+e_{2}}(t)\Bigl{]}-\frac{\overline{\lambda}\Delta t}{4(a_{0})^{2}}\prod_{i=1}^{2}\Bigl{[}h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl{]}+\sigma\sqrt{12\Delta t}R_{n}(t).$ (51)
Here, $\sigma^{2}=\frac{2D}{(a_{0})^{2}}$. The variable $R$ corresponds to uniformly distributed random numbers ranging from $-\frac{1}{2}$ to $\frac{1}{2}$. The prefactor $\sigma\sqrt{12\Delta t}$ ensures that the noise possesses an equivalent second moment as Gaussian noise integrated over a time interval $\Delta t$. Equation (51) can be conveniently expressed using dimensionless quantities $\tilde{h}=\frac{h}{h_{0}}$, $\tilde{t}=\frac{t}{t_{0}}$, and $\tilde{a}_{0}=\frac{a_{0}}{x_{0}}$, where the natural units are defined as $h_{0}=\frac{\nu}{\lambda}$, $t_{0}=\frac{\nu^{2}}{\sigma^{2}\lambda^{2}}$ and $x_{0}=\sqrt{\frac{\nu^{3}}{\sigma^{2}\lambda^{2}}}$. In these units Eq. (51) becomes
$\tilde{h}_{\bf n}(\tilde{t}+\Delta\tilde{t})$ $=\tilde{h}_{\bf n}(\tilde{t})+\frac{\Delta\tilde{t}}{(\tilde{a}_{0})^{2}}\sum_{i=1}^{2}\biggl{\{}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-2\tilde{h}_{\bf n}(\tilde{t})+\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}-\frac{1}{8}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}^{2}\biggl{\}}+\frac{\alpha\Delta\tilde{t}}{2(\tilde{a}_{0})^{2}}\Bigl{[}\tilde{h}_{\bf n+e_{1}+e_{2}}(\tilde{t})$
$-\tilde{h}_{\bf n+e_{1}-e_{2}}(\tilde{t})+\tilde{h}_{\bf n-e_{1}-e_{2}}(\tilde{t})-\tilde{h}_{\bf n-e_{1}+e_{2}}(\tilde{t})\Bigl{]}-\frac{\beta\Delta\tilde{t}}{4(\tilde{a}_{0})^{2}}\prod_{i=1}^{2}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}+\sqrt{12\Delta\tilde{t}}R_{n}(\tilde{t}).$ (52)
Figure 10: The variation of width ${\mathcal{W}}$ with time $t$ is presented on a log-log scale for different system sizes with two parameter sets: (a) $\alpha=0.50$, $\beta=2$, corresponding to the logarithmic rough region from RG analysis, with system sizes $L=100$, $300$, $500$, $800$, $1000$, and $1600$; (b) $\alpha=0.50$, $\beta=0.042$, corresponding to the algebraic rough region from our RG analysis, with system sizes $L=100$, $200$, $300$, $400$, and $500$. The red dashed line indicates a linear fit in the linear region on a log-log scale, yielding a slope of $\tilde{\beta}=0.234\pm 1.644\times 10^{-5}$ (see text).
For our simulations we chose $a_{0}=1$ and $\sigma=0.10$. First we discuss the numerical results for the logarithmic rough phase. For that, we set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$ and $\overline{\lambda}=4$ such that $|\alpha|<1$ and $|\beta|>1$, which corresponds to the logarithmic rough phase as predicted by our RG calculations (see main text). For a system of size $L$, we reached the saturation region (after the time after which width $\cal W$ saturates) by simulating Eq. (52) up to time $T\approx L^{2}$ in time step of $\Delta t=0.05$ as shown in Fig. 10(a) for different system sizes. Width $\cal W$ at a time $t$ was calculated using ${\cal W}=\sqrt{\langle\tilde{h}({\bf x},t)^{2}\rangle-\langle\tilde{h}({\bf x},t)\rangle^{2}}$. Figure 10(a) shows the variation of $\cal W$ with time $t$ in the growth and saturation region on a log-log scale and correspond to a single noise realization. The plot clearly shows that growth nearly saturates before reaching time $T\approx L^{2}$. To further investigate the saturation region, we extended the simulation for an additional $10^{6}$ time steps. During these $10^{6}$ time steps, we performed $N$ noise averages, with different choice of $N$ for different system sizes. For each of these time steps width ${\cal W}$ was calculated and averaged over $N$. The final saturated width ${\mathcal{W}_{\text{sat}}}$ was obtained by performing an average over $100000$ realizations of ${\cal W}_{\text{sat}}$ obtained after saturation for different system sizes $L=100$, $300$, $500$, $800$, $1000$ and $1600$. For $L=100$, $300$ noise average $N$ was set to $100$. For $L=500$ $N$ was set to $50$. And for $L=800$, $1000$ and $1600$ $N$ was set to $20$. Variation of ${\cal W}_{\text{sat}}$ with $\ln L$ in the saturation region is shown in the main text Fig. 3(a).
Next we discuss the numerical results for the algebraic rough phase. In this region, width ${\cal W}$ is expected to scale as $\mathcal{W}(L,t)\sim L^{\chi}f(\frac{t}{L^{z}})$ where, $f(x)$ is a dimensionless scaling function. With parameters $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$ and $\overline{\lambda}=1$ such that $|\alpha|<1$ and $|\beta|<1$, (which should correspond to the algebraic rough phase from RG analysis), we run the simulation for $L=100$, $200$, $300$, $400$ and $500$. For both of the growth and saturation region ${\cal W}$ was calculated at each time step $\Delta t=0.01$ with $N$ noise averages up to time $T$. For system size $L=100$, $200$, $300$, $400$ and $500$ the Eq. (52) was simulated up to time $T=5000$, $10000$, $15000$, $25000$ and $30000$ with time step $\Delta t=0.01$. $L=100$ to $400$ was noise averaged with $N=100$. For $L=500$, $N=50$ was taken. Refer to Fig. 10(b) for the variation of ${\cal W}$ over time on a log-log scale. The red dashed line indicates the slope of the growth region obtained through linear fitting, yielding the growth exponent $\tilde{\beta}=0.234\pm 1.644\times 10^{-5}$. Next we switch over to the saturation region of ${\cal W}$. With $T$ defined earlier for different $L$, we reached the saturation region of ${\cal W}$. See Fig. 10(b). Then we numerically solved the equation of motion (11) for further $100000$ time steps with $\Delta t=0.01$. And for each of these $100000$ time steps we have taken $N=500$ noise averages for all $L=100$ to $500$. For each of these time steps ${\cal W}$ was calculated and averaged over $N$. The final ${\mathcal{W}_{\text{sat}}}$ was calculated by averaging over 100000 realizations in the saturated regime. Variation of ${\cal W}_{\text{sat}}$ with $L$ in the saturation region is shown in Fig. 3(b) of the main text.

### v2

@@
 These are described by an anisotropic Kardar-Parisi-Zhang equation.
-Depending upon the anisotropy parameters, the membrane can be large-scale anisotropic and logarithmically rough with translational quasi long range order and orientational long range order, together with the relaxational dynamics being logarithmically faster than ordinary diffusion.
-For other choices of the anisotropy parameters, the membrane is either effectively isotropic and algebraically rough with translational short, but orientational long range order, or crumpled.
+Depending upon the anisotropy parameters, the membrane is either effectively isotropic and algebraically rough with translational short, but orientational long range order, or unstable, suggestive of membrane crumpling.
 The generality of the observations about the dynamical properties of the plasma membranes in live biological cells Alberts et al.
@@
 We consider an anisotropic membrane.
-For simplicity, we assume a fixed background for the membrane, i.e., the embedding fluid momentum is not conserved.
-We argue that the hydrodynamic equation for the undulation modes of such a membrane given by a single-valued height field $h({\bf x},t)$ is given by
-$\frac{\partial h}{\partial t}+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (1)
-Both $\lambda_{ij}$ and $\nu_{ij}$ are symmetric matrices. Noise $f$ has zero-mean and a variance given by
-$\langle f({\bf x},t)f(0,0)\rangle=2D\delta^{2}(x)\delta(t).$ (2)
+We argue that the, assuming Rouse dynamics Cai and Lubensky (1995) for simplicity, hydrodynamic equation for the undulation modes of an anisotropic membrane given by a single-valued height field $h({\bf x},t)$ is given by
+$\frac{\partial h}{\partial t}+\frac{\lambda}{2}{({\bm{\nabla}}h)^{2}}=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (1)
+Here $\nu_{ij}$ is symmetric under the interchange of $i$ and $j$, and is an effective anisotropic tension. Noise $f$ has zero-mean and a variance given by
+$\langle f({\bf x},t)f(0,0)\rangle=2D\delta^{2}({\bf x})\delta(t).$ (2)
 In the absence of any detailed balance, the noise strength $D$ has no specific relation to the damping coefficient matrix $\nu_{ij}$.
-For arbitrary $\nu_{ij}$ and $\lambda_{ij}$, Eq.
+For arbitrary $\nu_{ij}$, Eq.
 (1) is an anisotropic generalization of the well-known isotropic Kardar-Parisi-Zhang (KPZ) equation Barabási and Stanley (1995).
-For simplicity and to reduce the number of parameters, we further impose invariance of Eq.
-(11) separately under $(x,y)\rightarrow(\pm y,\pm x)$ in two dimensions (2D).
-We thus set $\nu_{xx}=\nu_{yy}=\nu>0$, $\nu_{xy}=\nu_{yx}=\overline{\nu}$, $\lambda_{xx}=\lambda_{yy}=\lambda$ and $\lambda_{xy}=\lambda_{yx}=\overline{\lambda}$, making it distinct from the one considered earlier Wolf (1991); Täuber and Frey (2002); Altman et al.
+Since in general $\nu_{ij}\neq 0$ for $i\neq j$, Eq.
+(1) is anisotropic.
+We set $\nu_{xx}=\nu_{x}$, $\nu_{yy}=\nu_{y}$, with $\nu_{x}\neq\nu_{y}$ both positive and $\nu_{xy}=\nu_{yx}=\overline{\nu}$ can be of any sign.
+Equation (1) is invariant under the joint transformation $(x,y)\to(-x,-y)$, which is of lower symmetry than the one considered earlier Wolf (1991); Täuber and Frey (2002); Altman et al.
 (2015).
-The most prominent result from direct numerical simulations (DNS) of Eq.
-(11) in two dimensions (2D) is that it admits membrane conformations $h({\bf x},t)$ having strikingly different statistical properties controlled by the model parameters.
-See Figs.
-1(a) and (b) and Figs.
-2(a) and (b) for representative snapshots of $h({\bf x},t)$ falling in three distinct phases (color bars represent $h$ relative to its mean) depending on the degree of the surface anisotropy controlled by $\overline{\nu},\overline{\lambda}$; see also the corresponding movies MOV1 [for Fig.
+See also Ref.
+da Silveira and Kardar (2003) for detailed studies on the stationary states of anisitropic KPZ equations.
+The most prominent result from direct numerical solutions (DNS) of Eq.
+(1) in 2D is that it admits different membrane conformations $h({\bf x},t)$ having strikingly distinct statistical properties controlled by the model parameters.
+See Fig.
+1 for representative snapshots of $h({\bf x},t)$ falling in two distinct phases (color bars represent $h$ relative to its mean) depending on the degree of the surface anisotropy controlled by $\nu_{x},\nu_{y},\overline{\nu}$.
+The corresponding movies MOV1 [for Fig.
 1(a)], MOV2 [for Fig.
-1(b)], MOV3(a) [for Fig.
-2(a)] and MOV3(b)[for Fig.
-2(b)] in Supplemental Material (SM) sup .
+1(c)] and MOV3 [for Fig.
+1(d)] are available in Supplemental Material (SM) sup .
 We characterize these phases by numerically solving Eq.
-(11) and measuring width ${\mathcal{W}}\equiv\sqrt{\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle}$, which as a function of time $t$ has an initial growing part, followed by a saturated part ${\cal W}_{\text{sat}}$ that fluctuates about a steady mean Barabási and Stanley (1995); $\overline{h}(t)$ is the mean height at time $t$.
-We find ${\mathcal{W}}_{\text{sat}}$ to scale as $(\ln L/a_{0})^{\tilde{\chi}}$ and $L^{\chi}$, respectively, for Fig.
-1(a) and Fig.
-1(b); $\tilde{\chi}\approx 0.43\pm 0.019,\chi=0.379\pm 0.008$ in our DNS studies, $L$ and $a_{0}$ are system size and small scale cutoff, respectively.
+(1) and measuring width ${\mathcal{W}}\equiv\sqrt{\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle}$, by averaging over different realisations, which as a function of time $t$ has an initial growing part, followed by a saturated part ${\cal W}_{\text{sat}}$ that fluctuates about a steady mean Barabási and Stanley (1995); $\overline{h}(t)$ is the spatially averaged mean height at time $t$.
+We find ${\mathcal{W}}_{\text{sat}}$ to scale as $L^{\chi}$, respectively, for Fig.
+1(a); $\chi=0.381\pm 0.002$; in our DNS studies, $L$ is system size.
 See Fig.
 3.
-Hence we name them logarithmic rough and algebraic rough phases respectively.
+Hence we name it algebraic rough phase.
 For Fig.
-2(a) and (b), no steady states are reached (within the computation time), giving ever increasing fluctuations, suggesting a complete loss of order, and hence we call it the crumpled phase.
-Snapshots in Figs.
-1(c) and (d) for respectively the logarithmic and algebraic rough phases for different sets of model parameters over enhanced ranges as used in Figs.
-2(a) and (b) for the crumpled phase indicate significantly larger fluctuations in the latter.
-See later for more details on DNS.
-Figure 1: Snapshots of the height profiles ($L=200$) in the steady states for (a) and (c) logarithmic rough phase, with parameters (a) $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$, and $\overline{\lambda}=4$, and (c) $\nu=0.60$, $\overline{\nu}=0.35$, $\lambda=5$, and $\overline{\lambda}=8$.
-Snapshots (b) and (d) correspond to the algebraic rough phase, with parameters (b) $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$, and $\overline{\lambda}=1$, and (d) $\nu=0.60$, $\overline{\nu}=0.35$, $\lambda=20$, and $\overline{\lambda}=2$.
-Notice the much larger fluctuations in the algebraic rough phase.
-See text.
-Figure 2: Snapshots of $h({\bf x},t)$ ($L=200$) in the crumpled phase with $\nu=1,\lambda=2$ showing much larger fluctuations than those in Fig.
-1.
-(a) $\overline{\nu}=2,\overline{\lambda}=2$ (with ripples at $\pi/4$), (b) $\overline{\nu}=-2,\overline{\lambda}=-2$ (with ripples at $43\pi/$ 4).
-See text.
-We show that the logarithmic rough phase is characterized by translational quasi-long-range order (QLRO) and orientational long-range order (LRO) with relaxation time scales being logarithmically faster than ordinary diffusion.
-Such a surface is effectively inversion-symmetric but has large-scale anisotropy.
-Further, the algebraically rough phase has effective isotropy with inversion-asymmetry and translational SRO and orientational LRO, akin to a 2D KPZ surface Kardar et al.
+1(c) and (d), no steady states are reached (within the computation time), giving ever increasing fluctuations, suggesting instability and a complete loss of order.
+Snapshot in Fig.
+1(b) represent the algebraic rough phase for a different set of model parameters over enhanced ranges as used in Figs.
+1(c) and (d) for the crumpled phase, indicate significantly larger fluctuations in the latter.
+See also later for the detailed numerical results from our DNS studies.
+Figure 1: Snapshots of the height profiles ($L=200$) in the steady states for (a) and (b) algebraic rough phase, with parameters (a) $\nu_{x}=0.50$, $\nu_{y}=0.45$, $\overline{\nu}=0.10$, $D=0.005$ and $\lambda=24$, and (b) $\nu_{x}=0.50$, $\nu_{y}=0.55$, $\overline{\nu}=0.20$, $D=0.005$ and $\lambda=24$.
+Snapshots (c) and (d) correspond to the crumpled phase, with parameters (c) $\nu_{x}=1.0$, $\nu_{y}=0.20$, $\overline{\nu}=1.0$, $\nu_{4}=0.10$, $D=1.0$ and $\lambda=2.20$, (with ripples at an arbitrary angle) and (d) $\nu_{x}=1.0$, $\nu_{y}=1.0$, $\overline{\nu}=2.0$, $\nu_{4}=0.10$, $D=1.0$ and $\lambda=2.30$ (with ripples at $\pi/4$).
+Notice the much smaller fluctuations in the algebraic rough phase, compared to the crumpled phase.
+(See text).
+We show that our model exhibits algebraic rough phase and is characterized by translational short-range order (SRO) and orientational long-range order (LRO) and specifically belongs to the KPZ universility class.
+Using DNS we verify for different sets of model parameters that the scaling exponents remain almost same indicating the putative algebraic rough phase is model parameter independent and indeed belongs to the KPZ universility class.
+Further, the algebraically rough phase is effectively isotropic with inversion-asymmetry and translational short range order (SRO) and orientational LRO, akin to a 2D KPZ surface Kardar et al.
 (1986); Tang et al.
-(1990); Barabási and Stanley (1995) and crumpled phase with translational and orientational SRO, respectively.
-These results should be contrasted with the corresponding isotropic case, which only has a rough phase.
+(1990); Barabási and Stanley (1995) and the crumpled phase has translational and orientational SRO, respectively.
 We now derive Eq.
 (1) and use it to establish the above results and obtain the phase diagrams.
+The membrane configurations are given by three-dimensional position vector ${\bf R}({\vec{s}},t)$, where ${\vec{s}}=(s_{1},s_{2})$ is a two-dimensional position vector embedded on the membrane, which label points on the membrane.
+The local membrane velocity ${\bf v}({\vec{s}},t)\equiv\partial{\bf R}/\partial t$, for a convected or embedded coordinate system Aris (1989); Hu et al.
+(2007).
+We now specialize for nearly flat inversion-asymmetric active membrane patches without any overhangs, for which the Monge gauge Nelson et al.
+(1989) is a convenient choice, in which $(s_{1},s_{2})\equiv(x,y)$ and ${\bf R}({\bf x})=({\bf x},h({\bf x})),\,{\bf x}=(x,y)$.
 The only hydrodynamic variable is $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al.
 (1989).
-We consider a nearly flat membrane patch to span at $z=0$ along the XY plane.
-Assuming simple relaxational dynamics, the general equation of motion of $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al.
+Due to the breakdown of inversion symmetry, the patch, considered permeable, is generically moving in a direction local normal to itself.
+A change in the membrane conformation $h({\bf x},t)$ can take place by the drift of the membrane due to activity and relaxational dynamics, which we assume to be Rouse dynamics for simplicity.
+The general equation of motion of $h({\bf x},t)$, measured with respect to an arbitrary base plane in the Monge gauge Nelson et al.
 (1989), reads
-$\frac{\partial h}{\partial t}=-\frac{\delta{\cal F}}{\delta h}+v_{\text{perm}}+f.$ (3)
-Here, ${\cal F}$ is a free energy of an equilibrium fluid membrane, assumed tensionless, and $v_{\text{perm}}$ is the permeation velocity locally normal to the plane of the membrane.
-Here, ${\cal F}=\int d^{2}x[\frac{K}{2}(\nabla^{2}h)^{2}+C\nabla^{2}h]$, ignoring anharmonic contributions, is a free energy functional, which governs the relaxation of the membrane to equilibrium in the absence of any activity; here, $K>0$ is the bending stiffness and ${\bm{\nabla}}\equiv(\partial/\partial x,\,\partial/\partial y)$ is the 2D gradient operator; $C$ is the spontaneous curvature that can be generically present in an inversion-asymmetric membrane and makes the membrane curve one way or the other depending upon its sign.
-In the Monge gauge approximation, valid for a planar membrane patch, the spontaneous curvature term being a total derivative term vanishes and makes no contribution to $\cal F$.
-Consider a nearly flat patch of an inversion-asymmetric, anisotropic fluid membrane placed in an isotropic, active suspension of actin filaments Marchetti et al.
+$\frac{\partial h}{\partial t}=v\sqrt{1+({\bm{\nabla}}h)^{2}}-\Gamma\frac{\delta F}{\delta h}+f,$ (3)
+[…]
+$v(h)\equiv v_{0}+\nu_{ij}\partial_{i}\partial_{j}h.$ (4)
+In the cellular context, $v_{0}$ sets the scale of the drift speed of the membrane, e.g., due to actin filament polymerization Sykes et al.
+(2010) (see also below).
+The second term on the rhs of (4) gives the contribution from the membrane conformation fluctuations to the drift.
+Assuming small fluctuations, we expand the square root in (3) and obtain (1).
+A possible microscopic realization of this would be a nearly flat patch of an inversion-asymmetric, anisotropic fluid membrane placed in an isotropic, active suspension of actin filaments Marchetti et al.
 (2013) grafted normally to the membrane locally.
-Let $\bf p$ be a three-dimensional unit vector ($p^{2}=1$) that gives the local orientation of the actin filaments and $\bf n$ be the local unit normal to the membrane; ${\bf n}=(\hat{z},-\nabla_{i}h)$ to the linear order in height fluctuations, in the present geometry with $i=x,y$.
-Locally normal grafting of the filaments implies ${\bf p}\cdot{\bf n}=1$.
-This yields $p_{j}=-\partial_{j}h\,(j=x,y)$ to the linear order in height fluctuations Sarkar and Basu (2013); Banerjee and Basu (2018).
-Now $v_{\text{perm}}$ Banerjee and Basu (2018), locally normal to an anisotropic membrane Barabási and Stanley (1995), reads
-$v_{\text{perm}}=(v_{0}+\xi_{ij}\partial_{i}p_{j}+\tilde{\xi}_{ij}p_{i}p_{j})\sqrt{1+({\bm{\nabla}}h)^{2}}$ (4)
-to the lowest order in small fluctuations $p_{i}$; $i=x,y$.
-In the cellular context, $v_{0}$ sets the scale of the drift speed of the membrane due to actin filament polymerization Sykes et al.
-(2010).The remaining terms in (4) give the drift due to local orientation fluctuations.
-Coefficient matrices $\xi_{ij}$ and $\tilde{\xi}_{ij}$ encode the anisotropy, which may originate from the anisotropy in the actomyosin cortex Mayer et al.
+Let $\bf p$ be a three-dimensional unit vector ($p^{2}=1$) that gives the local orientation of the actin filaments and $\bf n$ be the local unit normal to the membrane; ${\bf n}=(\hat{z},-\nabla_{i}h)$ to the linear order in height fluctuations ($i=x,y$).
+Locally normal grafting of the filaments implies ${\bf p}\cdot{\bf n}=1$, giving $p_{i}=-\partial_{i}h$ ($i=x,y$) and $p_{z}=1$ to the linear order in fluctuations.
+The actin filaments exert forces locally normally to the membrane Chen and Mikhailov (2010), giving a three-dimensional local permeation flow velocity
+${\bf v}\sim v_{0}{\bf p}+{\bm{\nabla}}\cdot[{\bm{\zeta}}{\bf p}{\bf p}].$ (5)
+Now, extracting the component of $\bf v$} along the membrane’s local normal and to linear order in $h$ yields $v$ as above Maitra et al.
+(2014); Banerjee and Basu (2018).
+Writing the second term on the rhs of (5) as the divergence of an active stress ${\bm{\sigma}}^{a}\equiv{\bm{\zeta}}{\bf p}{\bf p}$ i.e.
+$\sigma_{ij}^{a}\equiv\zeta_{ij\alpha\beta}p_{\alpha}p_{\beta}$, we see that $\nu_{ij}$ is proportional to $\sigma_{ij}^{a}$.
+Matrix ${\bm{\zeta}}\equiv\zeta_{ij\alpha\beta}$ encodes the anisotropy in the membrane, e.g., anisotropy in the actomyosin cortex Mayer et al.
 (2010), or anisotropic inclusions in the membrane Fošnarič et al.
 (2005).
-Expanding and substituting for $p_{i}$ and retaining only the lowest order terms in gradients and nonlinearities, we obtain Eq.
-(11), where $\lambda_{ij}=\tilde{\xi}_{ij}$ is a measure of the average membrane speed due to fluctuations, and $\nu_{ij}=-\xi_{ij}$ is an anisotropic tension.
-Structures of these anisotropy parameters should depend explicitly on the specific model representation.
-Figure 3: Log-log plots of (a) ${\mathcal{W}_{\text{sat}}}$ versus $\ln L$ ($a_{0}=1$) for $\alpha=0.50$, $\beta=2$ (logarithmic rough phase) and $L=100$, $300$, $500$, $800$, $1000$ and $1600$.
-Red dashed line represents linear fit to the numerical results displayed on a log-log scale with a slope $\tilde{\chi}=0.43\pm 0.019$.
-(b) ${\mathcal{W}_{\text{sat}}}$ with $L$ for $\alpha=0.50$, $\beta=0.042$ (algebraic rough phase) and $L=100$, $200$, $300$, $400$ and $500$.
-Red dashed line represents linear fit to the numerical results displayed on a log-log scale with a slope $\chi=0.379\pm 0.008$ (see text).
+The forms of these anisotropy parameters should depend explicitly on the specific model representation.
+In the absence of anisotropy, in which case the strength of the active stress is described by a scalar parameter, say $\Delta\mu$, a positive (negative) $\Delta\mu$ refers to extensile (contractile) activity Marchetti et al.
+(2013).
+Here, however the active stress is characterized by three parameters, and hence a simple extensile/contractile characterization is no longer possible.
+Figure 2: (a) Two distinct phases in $\beta$-$\alpha$ plane.
+White region is the KPZ like putative algebraic rough phase.
+(b) The RG flow diagram in the $\beta$-$\alpha$-$g$ space.
+Arrows pointing away from the $\beta$-$\alpha$ plane represent the direction of RG flows, indicating a strong coupling algebraic rough phase.
+Green regions in (a) and (b) correspond to crumpled phase (see text).
 We now theoretically explain the above numerical observations, quantitatively characterizing the phases, by using Eq.
 (1).
-Equation (11) admits a pseudo-Galilean invariance under the transformation $t^{\prime}=t,\,x_{i}^{\prime}=x_{i}+\lambda_{ij}u_{j}t,\,h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+u_{i}x_{i}+\frac{\lambda_{ij}}{2}u_{i}u_{j}t$ sup .
+Equation (1) admits a pseudo-Galilean invariance under the transformation $t^{\prime}=t,\,x_{i}^{\prime}=x_{i}+\lambda u_{i}t,\,h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+u_{i}x_{i}+\frac{\lambda}{2}u^{2}\,t$ sup .
 In (1), $\nu_{ij}$ plays the role of an (active) anisotropic surface tension.
-With $\nu>0$, linear stability requires that $\nu\pm\overline{\nu}\geq 0$, or $|\alpha|\leq 1$, i.e.
-$-1\leq\alpha\leq 1$; otherwise instability ensues.
-This means $\alpha_{c}=\pm 1$ is a threshold of $\alpha$, with larger $|\alpha|$ implies instability.
-This is the anisotropy-induced crumpling mentioned earlier in this Letter, which can occur even at zero noises.
-This instability threshold is independent of wavevector $q$, any ensuing crumpling actually takes place at all scales, meaning an arbitrarily small membrane will be crumpled, provided $|\alpha|>1$.
-This is reminiscent of the strongly crumpled phase in tensionless equilibrium asymmetric tethered membranes Banerjee et al.
+With $\nu_{x},\nu_{y}>0$, linear stability requires that $\nu_{x}\nu_{y}\geq\overline{\nu}^{2}$, i.e.
+$\beta\geq\alpha^{2}$; otherwise instability ensues.
+Here $\beta=\frac{\nu_{y}}{\nu_{x}}$ and $\alpha=\frac{\overline{\nu}}{\nu_{x}}$.
+If $\alpha=0$, Eq.
+(1) is invariant under (separately) $x\rightarrow-x$ or $y\rightarrow-y$; for $\alpha\neq 0$, the invariance is only under the joint inversion of $x$ and $y$.
+This means $\alpha_{c}=\pm\sqrt{\beta}$ is a threshold of $\alpha$, with larger $|\alpha|$ implies instability.
+This is the anisotropy-induced instability mentioned earlier in this Letter, which can occur even at zero noises, provided $|\alpha|>\sqrt{\beta}$.
+This then means in general the variance of the local height fluctuations $\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle$ and local normal fluctuation $\delta{\bf n}=-{\bm{\nabla}}h$, $\langle(\delta{\bf n}({\bf x},t))^{2}\rangle$ diverge.
+Can this instability be suppressed by including stabilizing higher order damping terms, e.g., $\nu_{4}\nabla^{4}h,\,\nu_{4}>0$ in (1)?
+Such a term can be generically present, but is irrelevant if there is no linear instability.
+In the event of linear instability, these higher order (in gradient) stabilizing terms are important.
+Their presence would ensure that at high enough wavevectors the system is linearly stable: If $\nu_{1},\nu_{2}$ are the eigenvalues of of $\nu_{ij}$, then for membranes of linear size $L<(>)L_{c}\sim\sqrt{\nu_{4}/|\text{min}(\nu_{1},\nu_{2})|}$ are dominated by the stabilizing hyperdiffusion (destabilizing diffusion) terms.
+In this case, $\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle$ and $\delta{\bf n}=-{\bm{\nabla}}h$, $\langle(\delta{\bf n}({\bf x},t))^{2}\rangle$ diverge not at every scale, but only if the linear system size $L$ exceeds a threshold that scales with $\sqrt{\nu_{4}}$ Mukherjee and Basu (2024).
+Such divergences of $\langle(\delta{\bf n}({\bf x},t))^{2}\rangle$ for sufficiently larger membranes indicates that such membranes with $L>L_{c}$ exceeding the instability threshold may be crumpled, although (1) cannot be used to study the crumpled phase; see also later and Ref.
+Mukherjee and Basu (2022).
+This instability is reminiscent of the strongly crumpled phase in tensionless equilibrium asymmetric tethered membranes Banerjee et al.
 (2019a, b).
-In the crumpled phase, the variance of the local normal fluctuations $\sim\langle({\bm{\nabla}}h)^{2}\rangle$ diverges for a sufficiently large but finite $L$, giving translational and orientational SRO.
-In this phase, the growth rate is maximum at an angle $\pi/4$ and $43\pi/$ 4, respectively, with the $x$-axis for $\alpha>1$ and $\alpha<-1$.
-This is reflected in the ripples observed Fig.
-2(a) and (b), and the corresponding movies MOV3(a) and MOV3(b) in SM sup .
-We now turn to the uncrumpled case with $|\alpha|<1$, and consider its stability and fluctuation properties.
-In the linear theory with $|\alpha|<1$, width ${\mathcal{W}}_{\text{sat}}\sim\sqrt{\ln(L/a_{0})}$.As in the usual KPZ equation Kardar et al.
+In this unstable phase, the variance of the local normal fluctuations $\sim\langle({\bm{\nabla}}h)^{2}\rangle$ diverges for a sufficiently large but finite $L$, giving translational and orientational SRO.
+In this phase, the growth rate is maximum forming ripples at a particular angle depending on parameters $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$; see Fig.
+1(c).
+Consider the special case where $\nu_{x}=\nu_{y}$ and $|\alpha|>1$ then ripples will form at an angle $\pi/4$ ($3\pi/4$) with the $x$-axis for $\alpha>1$ ($\alpha<-1$); see SM sup .
+See the ripples in Fig.
+1(d), forming at an angle $\pi/4$, and the corresponding movies MOV2 and MOV3 in SM sup .
+We now turn to the stable case with $|\alpha|<\sqrt{\beta}$, and consider its stability and fluctuation properties.
+In the linear theory with $|\alpha|<\sqrt{\beta}$, width ${\mathcal{W}}_{\text{sat}}\sim\sqrt{\ln(L/a_{0})}$.
+As in the usual KPZ equation Kardar et al.
 (1986), the combination of noises and nonlinearities here can change the linear theory results substantially.
 To study the nonlinear effects in the present model (1), we perform a one-loop perturbative dynamic renormalization group (RG) analysis of Eq.
-(11) with $|\alpha|<1$.
+(1) with $|\alpha|<\sqrt{\beta}$.
 The Wilson dynamic RG method Täuber (2014); Forster et al.
@@
 (1986); Barabási and Stanley (1995); Täuber (2014).
-Galilean invariance of (11) ensures that $\lambda_{ij}$ is unrenormalized, which is an exact statement and the analogue of the nonrenormalization of the nonlinear coupling constant in the usual KPZ equation Barabási and Stanley (1995); Frey and Täuber (1994).
-However, there are diverging one-loop corrections to $\nu,\,\overline{\nu}$ and $D$, which are represented by the one-loop Feynman diagrams (see SM) sup .
-Dimensional analysis allows us to identify an effective dimensionless coupling constant $g$ having a critical dimension 2 and a dimensionless ratio $\beta$ defined as
-$g=\frac{\lambda^{2}D}{8\pi(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}},\;\beta=\frac{\overline{\lambda}}{\lambda},$ (5)
-along with $\alpha$ defined earlier. The RG differential recursion relations ($\exp(l)$ is a dimensionless length) are
-$\frac{dD}{dl}=D\bigg{[}z-2-2\chi+g\{\frac{(2+\alpha^{2})+\beta^{2}(1+2\alpha^{2})-6\alpha\beta}{2(1-\alpha^{2})}\}\bigg{]},$ (6)
-$\frac{d\nu}{dl}=\nu[z-2-g\{\frac{\alpha\beta}{2}-\frac{\beta^{2}}{2}\}],$ (7)
-$\frac{d\overline{\nu}}{dl}=\overline{\nu}[z-2-g\{\frac{1}{2}-\frac{\beta}{2\alpha}\}],$ (8)
-$\frac{dg}{dl}=-g^{2}(\beta^{2}-1),$ (9)
-together with $d\beta/dl=d/dl(\overline{\lambda}/\lambda)=0$.
-Here, $z$ and $\chi$ are the dynamic and roughness exponents, respectively Barabási and Stanley (1995)In the linear theory, i.e., with $\lambda_{ij}=0$, $z=2,\,\chi=0$, giving translational QLRO.
-Equation (9) gives that if $\beta^{2}>1$, then $g(l)\sim\frac{1}{(\beta^{2}-1)l}$ for large “RG time” $l$.
-On the other hand, if $\beta^{2}<1$, $g(l)$ diverges as $l\rightarrow l_{c}\sim{\cal O}(1)$ from below, where $l_{c}$ is a nonuniversal model parameter-dependent length.
-Of course, we cannot follow all the way to $g(l)$ diverging; as soon as $g(l)$ becomes ${\cal O}(1)$, the perturbation theory breaks down.
-This is the strong coupling algebraic rough phase or the model, whose scaling properties are not amenable to perturbative RG treatments, and is analogous to the strong coupling rough phase in the 2D KPZ equation Barabási and Stanley (1995); Tang et al.
+Galilean invariance of (1) ensures that $\lambda$ is unrenormalized, which is an exact statement and the analogue of the nonrenormalization of the nonlinear coupling in the usual KPZ equation Barabási and Stanley (1995); Frey and Täuber (1994); see also SM sup .
+However, there are diverging one-loop corrections to $\nu_{x},\,\,\nu_{y},\,\overline{\nu}$ and $D$, which are represented by the one-loop Feynman diagrams (see SM) sup .
+Dimensional analysis allows us to define an effective dimensionless coupling constant $g$ having a critical dimension 2 as
+$g=\frac{\lambda^{2}D}{16\pi(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}},$ (6)
+along with $\alpha$, $\beta$ defined earlier. The RG differential recursion relations ($\exp(l)$ is a dimensionless length) are
+$\frac{dD}{dl}=D\bigg[z-2-2\chi+g\frac{3(1+\beta^{2})/4+\beta/2+\alpha^{2}}{(\beta-\alpha^{2})}\bigg],$ (7)
+$\frac{d\nu_{x}}{dl}=\nu_{x}[z-2-g(1/2-\beta/2)],$ (8)
+$\frac{d\nu_{y}}{dl}=\nu_{y}[z-2-g(1/2-1/(2\beta))],$ (9)
+$\frac{d\overline{\nu}}{dl}=\overline{\nu}[z-2-g],$ (10)
+$\frac{dg}{dl}=2g^{2},$ (11)
+$\frac{d\alpha}{dl}=-g(\alpha+\alpha\beta)/2,$ (12)
+$\frac{d\beta}{dl}=g(1-\beta^{2})/2.$ (13)
+Here, $z$ and $\chi$ are the dynamic and roughness exponents, respectively Barabási and Stanley (1995).
+Equation (11) shows that $g(l)$ diverges as $l\rightarrow l_{c}\sim{\cal O}(1)$ from below, where $l_{c}$ is a nonuniversal model parameter-dependent length.
+Of course, we cannot follow all the way to $g(l)$ diverging; as soon as $g(l)$ becomes ${\cal O}(1)$, the perturbation theory breaks down and hints a strong coupling algebraic rough phase similar to the usual 2D KPZ equation Tang et al.
+(1990); Barabási and Stanley (1995).
+Assuming $g=g^{*}$ at the putative perturbatively inaccessible strong coupling fixed point (SCFP), the flow equations (12) and (13) for $\alpha$ and $\beta$ respectively yield $\beta=1$ and $\alpha=0$ as the stable fixed points.
+We therefore speculate that, in the long-wavelength limit ($l\to\infty$), $\nu_{x}\to\nu_{y}$ and $\overline{\nu}\to 0$, indicating that the system becomes effectively isotropic and belongs to the KPZ universality class.
+Figure 3: Log-log plots of $\mathcal{W}$ versus $t$ (a) for $L=500$ with $\nu_{x}=0.50$, $\overline{\nu}=0.10$, $D=0.005$, $\lambda=24$, and different values of $\nu_{y}$.
+Inset: linear-scale plot for $L=300$.
+(b) For $L=500$ with $\nu_{y}=0.55$ and varying $\overline{\nu}$ (other parameters as in (a)).
+Inset: plot for $L=300$ in linear-scale.
+(c) For $\nu_{y}=0.55$, $\overline{\nu}=0.20$, (other parameters as in (a)) and system sizes $L=100$ to $500$.
+The red dashed line indicates a linear fit with slope $\tilde{\beta}=0.22\pm 2.6\times 10^{-5}$.
+(d) Log-log plot of Saturation width $\mathcal{W}_{\text{sat}}$ versus $L$ for the same parameters as in (c), with a linear fit yielding $\chi=0.38\pm 0.002$ (see text).
+Figure 2(a) shows the phase diagram in the $\beta$-$\alpha$ plane.
+The green regions at the top and bottom, characterized by $|\alpha|>\sqrt{\beta}$, correspond to the crumpled phase driven by a linear instability.
+The intervening white region, defined by $|\alpha|<\sqrt{\beta}$, is linearly stable but perturbatively inaccessible within the RG framework.
+Figure 2(b) displays the corresponding RG flow lines in the full $\beta$-$\alpha$-$g$ space.
+Flow lines originating within the whitish paraboloid region (bounded by $|\alpha|<\sqrt{\beta}$) escape to a SCFP that is perturbatively inaccessible, reminiscent of the behavior in the 2D isotropic KPZ equation Barabási and Stanley (1995); Tang et al.
+(1990)
+To complement the predictions of this RG analysis, we perform DNS of Eq.
 (1).
-The RG recursion relation of $\alpha$ reads
-$\frac{d\alpha}{dl}=\frac{g}{2}\bigr{[}\alpha^{2}\beta-\alpha\beta^{2}-\alpha+\beta\bigr{]}.$
-Setting $\frac{d\alpha}{dl}=0$, we obtain
-$\alpha^{*}=1/\beta,\beta$
-as the fixed points of $\alpha$.
-Thus in the perturbatively accessible region $\beta^{2}>1$, $\alpha^{*}=1/\beta$, $(\beta)$ is stable (unstable) fixed line.
-Focusing on the stable case $\beta^{2}>1$ and using the linear theory results $z=2$, $\chi=0$ we find $\nu(l)\sim\nu_{0}\sqrt{l},\,\overline{\nu}(l)\sim\overline{\nu}_{0}\sqrt{l},\,D(l)\sim D_{0}\sqrt{l}$ Mukherjee and Basu (2024), with $\nu_{0},\overline{\nu}_{0},D_{0}$ being the “initial” values of $\nu(l),\,\overline{\nu}(l),D(l)$ respectively.
-This gives logarithmically faster relaxation than ordinary diffusion, time-scale $t\sim r^{2}/\sqrt{\ln(r/a_{0})}$, where $r$ is a distance, meaning logarithmic breakdown of dynamic scaling.
-Similar logarithmically modulated noise strength and damping are also found in Refs.
-Haldar et al.
-(2023a, b); Jana et al.
-(2024).
-We further get
-${\mathcal{W}}_{\text{sat}}\sim(D_{0}/{\mathcal{A}})\sqrt{\ln{L}/{a_{0}}},$ (10)
-at the stable RG fixed point, where ${\mathcal{A}}\equiv[{4\pi\nu_{0}({1-(\frac{\overline{\nu}_{0}}{\nu_{0}})^{2}})^{1/2}}]^{1/2}$.
-This corresponds to a logarithmically rough surface with translational QLRO, statistically identical to a 2D Edward-Wilkinson (EW) surface Barabási and Stanley (1995); Edwards and Wilkinson (1982).
-An effective temperature $T_{\text{eff}}$ defined via $T_{\text{eff}}=D(l)/\nu(l)=D_{0}/\nu_{0}$, i.e., unchanged from its microscopic value.
-Nonetheless, due to its logarithmically modulated noise and damping, at this RG fixed point our model is different from the 2D EW equation.
-While the equal-time height-difference correlator is same as in the EW equation, the more general time-displaced correlator is different.
-With $\alpha^{*}=1/\beta$ and $\beta^{2}>1$ at the RG fixed point, the logarithmic rough phase is macroscopically anisotropic.
-In the perturbatively inaccessible case with $\beta^{2}<1$, we have $\alpha^{*}=\beta$ as the stable fixed point, implying large-scale isotropy.
-In this case $d\nu/dl,d\overline{\nu}/dl=0,dD/dl>0$, reminiscent of the RG flow equations in the 2D isotropic KPZ equation Barabási and Stanley (1995).
-Figure 4: (a) Different phases and the RG flow lines of $\alpha$ in $\beta$-$\alpha$ plane.
-Yellow regions are the logarithmic rough phase.
-Central white region is the KPZ like putative algebraic rough phase.
-The black lines in the rough phase show the locus of the fixed point value $\alpha^{*}$ as a function of $\beta$.
-(b) The RG flow diagram in the $\beta$-$\alpha$-$g$ space.
-Arrows pointing toward (away from) the $\beta$-$\alpha$ plane represent the direction of RG flows, indicating stable (unstable) regions.
-Green regions in (a) and (b) correspond to crumpled phase (see text).
-Figure 4(a) shows the phases and the RG flow lines of $\alpha$ in $\beta$-$\alpha$ plane.
-Since $d\beta/dl=0$, the RG flow lines in the $\beta$-$\alpha$ plane are perpendicular to the $\beta$-axis.
-The green horizontal regions at the top and bottom of Fig.
-4(a) have $|\alpha|>1$, corresponding to the crumpled phase.
-The remaining region of this figure with $|\alpha|<1$, corresponding to linear stability, has two parts: The regions with $\beta^{2}>1$ (marked yellow) are perturbatively accessible logarithmically rough phase, whereas the region with $\beta^{2}<1$ is perturbatively inaccessible.
-Figure 4(b) shows the RG flow lines in the $\beta$-$\alpha$-$g$ space.
-Flow lines with initial values of $g$ in the yellow region of $\alpha$-$\beta$ plane flow to zero giving logarithmic roughness.
-Flow lines with initial values of $g$ inside the whitish cuboid region flow out to the perturbatively inaccessible strong coupling fixed point indicating KPZ-like algebraic roughness.
-In our DNS studies, used to corroborate our RG results and characterize the phases, we discretize Eq.
-(11) by standard forward-backward differences on a square grid with lattice spacing $a_{0}=1$, and numerically solve it by using the Euler algorithm; see SM sup for more details.
-We calculate ${\mathcal{W}}$ for model parameters falling in the (i) yellow, stable region (logarithmically rough), (ii) unstable, white region (speculated to be algebraically rough) and (iii) green, linearly unstable (crumpled) region in the phase diagram 4(a).
-For the logarithmic rough phase, we have set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$ and $\overline{\lambda}=4$ such that $|\alpha|<1$ and $|\beta|>1$ and system sizes $L=100$, $300$, $500$, $800$, $1000$ and $1600$ and time step $\Delta t=0.05$ in our simulations.
-Numerical measurements of ${\cal W}_{\text{sat}}$ for various $L$ give $\tilde{\chi}\approx 0.43\pm 0.019$, close to the analytically obtained value $\tilde{\chi}=0.50$, firmly establishing the logarithmic rough phase.
-Next we set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$ and $\overline{\lambda}=1$ such that $|\alpha|<1$ and $|\beta|<1$, corresponding to $dg(l)/dl>0$ [see Eq.
+We discretize the equation using standard forward-backward finite differences on a square grid with lattice spacing $a_{0}=1$, and integrate it numerically via the Euler algorithm with a fixed time step $\Delta t=0.01$; further implementation details are provided in the SM sup .
+We compute the interface width $\mathcal{W}$} for parameter sets lying in both the (i) speculated algebraically rough regime and (ii) linearly unstable (crumpled) regime of Fig.
+2(a).
+Now we focus on the DNS results within the speculated algebraic rough regime, which is argued to be statistically identical to the isotropic 2D KPZ equation.
+We chose parameter space such that $dg(l)/dl>0$ [see Eq.
 (11)].
-Time step $\Delta t=0.01$.
-With ${\cal W}\sim t^{\tilde{\beta}}$ initially, we obtained $\tilde{\beta}\approx 0.234\pm 1.644\times 10^{-5}$; see also SM sup .
-Furthermore, numerically measuring ${\mathcal{W}}_{\text{sat}}$ for $L=100$, $200$, $300$, $400$ and $500$ give $\chi\approx 0.379\pm 0.008$, and hence $z=\chi/\tilde{\beta}=1.619$ with $\chi+z\approx 2$ (as it should be).
+Fig.
+3(a) shows the $\mathcal{W}(t)$ for various $\nu_{y}$ values with fixed $\nu_{x},\overline{\nu}$.
+The linear growth regime exhibits an approximately constant slope across all cases, suggesting that the growth exponent $\tilde{\beta}$ is insensitive to $\nu_{y}$.
+However, the $\mathcal{W}_{\text{sat}}$ increases with decreasing $\nu_{y}$, as lower $\nu_{y}$ (for a fixed $\nu_{x}$) drives the system closer to the instability threshold.
+See Fig.
+2(a).
+Fig.
+3(b) presents $\mathcal{W}(t)$ for different $\overline{\nu}$.
+Again, the growth exponent remains nearly unchanged, while $\mathcal{W}_{\text{sat}}$ increases with $|\overline{\nu}|$, consistent with approaching the unstable region at fixed $\nu_{x}$.
+See Fig.
+2(a).
+Fig.
+3(c) shows variation of $\cal W$} over time for different $L$.
+The red dashed line indicates the slope of the growth region obtained through linear fitting, yielding $\tilde{\beta}=0.22\pm 2.6\times 10^{-5}$.
+Furthermore, numerically measuring ${\mathcal{W}}_{\text{sat}}$ for $L=100$, $200$, $300$, $400$ and $500$ give $\chi\approx 0.38\pm 0.002$, and hence $z=\chi/\tilde{\beta}=1.72$ with $\chi+z\approx 2$ (as it should be).
+See Fig.
+3(d).
 Our results on $\chi$ confirms an algebraically rough phase with translational SRO and orientational LRO, as speculated from our RG calculations above.
-The value of $\chi=0.379$ shows reasonable agreement with the existing results Oliveira et al.
+The value of $\chi=0.38$ compares reasonably with the existing results Oliveira et al.
 (2013); Aarão Reis (2004); Kondev et al.
@@
 (2017); Ódor et al.
-(2009); Halpin-Healy (2012); Chin and den Nijs (1999); Colaiori and Moore (2001); Gomes-Filho et al.
-(2021) and also with a mode coupling theory prediction Bhattacharjee (1998), indicating scaling behavior statistically same as in the strong coupling rough phase of the 2D KPZ equation.
-These DNS results lend further credence to a strong coupling rough phase with large-scale isotropy.
+(2009); Halpin-Healy (2012); Chin and den Nijs (1999); Colaiori and Moore (2001); Halpin-Healy (2013) and also not far from a mode coupling theory Bhattacharjee (1998) and functional RG Halpin-Healy (1990) predictions, indicating 2D KPZ-like scaling behavior lending further credence to a strong coupling rough phase belonging to the 2D KPZ universality class.
+To explore the unstable region, we introduce a regularizing term $\nu_{4}\nabla^{4}h$ to Eq.
+(1).
+For DNS in the unstable region (i.e green region in Fig.
+2(a), (b)) we choose $\nu_{4}=0.10$, $\Delta t=0.01$.
+However, the interface width still diverges rapidly after a short duration, possibly indicating that additional higher-order nonlinear terms, neglected in (1) by using scaling arguments, are now required to describe this region reliably.
+Fig.
+1(c) [(d)], with $L_{c}=0.46$ $L_{c}=0.32$], both much smaller than the membrane size $L=200$, display the height snapshots just before divergence occurs.
 So far we have considered a nearly flat patch within a Monge parametrization, which holds up to length scales at which curvature $C$ becomes important.
-Within this scale, while microscopically the membrane is inversion-asymmetric, the effective long wavelength behavior is symmetric in the logarithmic rough phase.
-The nonlinear terms with coefficients $\lambda_{ij}$ break the inversion symmetry.
-However, coupling $g$ vanishes in the long wavelength limit (albeit slowly enough to introduce logarithmic modulations to the noise and damping).
-At larger scales, when $C$ should be important, the spontaneous curvature term scales as $C/r$, whereas the surface tension term scales as $r^{0}$, making $C$ increasingly unimportant (relative to the surface tension) at larger length scales.
-We thus expect the membrane to remain effectively inversion-symmetric (and statistically flat) at arbitrarily large scales.
-In the algebraic rough phase, $g$ is not expected to vanish, giving inversion-asymmetry to survive in the long wavelength limit.
-Nonetheless, with expected $z<2$ the surface tension continues to dominate over $C$, ultimately giving a statistically flat but inversion-asymmetric membrane.
-At the onset of the crumpled phase, the effective surface tension vanishes, making $C$ important.
-This is beyond the scope of the present study.
-We have thus shown that the hydrodynamic theory for a moving active anisotropic inversion-asymmetric fluid membrane is given by an anisotropic KPZ equation, giving a logarithmically rough phase an algebraically rough phase and a crumpled phase.
-The average membrane velocity $\lambda_{ij}\langle{\nabla_{i}h}\nabla_{j}h\rangle$, an observable feature and proportional to $g$, vanishes in the inversion-symmetric logarithmic rough phase, but is finite in the asymmetric algebraically rough phase.
-It would be interesting to explore connections between our predicted crumpled phase and observed reduction of membrane tension to even negative values during membrane fusion Solon et al.
+At larger scales, when $C$ could be important, the spontaneous curvature term scales as $C/r$.
+In the algebraic rough phase, $g$ should not vanish, with inversion-asymmetry surviving in the long wavelength limit.
+Nonetheless, with expected $z<2$, $\nu_{ij}$ continues to dominate over $C$ (in a scaling sense), ultimately giving a statistically flat but inversion-asymmetric membrane.
+For the crumpled phase, it is not possible to comment on the large-scale structure of the membrane.
+We have thus shown that the hydrodynamic theory for a moving active anisotropic fluid membrane is given by an anisotropic KPZ equation, giving an algebraically rough phase, that in spite of the anisotropy belongs to the isotropic 2D KPZ universality class (and hence isotropy is en emergent symmetry of the model in the long wavelength limit) and an unstable phase that is speculated to be a crumpled phase.
+Existence of this putative crumpled phase is linked to sufficiently strong breaking of the inversion symmetry of the model.
+Our results show that the fluctuation-induced average membrane velocity $\frac{\lambda}{2}\langle(\nabla h)^{2}\rangle$, proportional to $g$ is finite in the algebraically rough phase.
+We are unable to comment on it in the unstable phase.
+In living cells, lamellipodia where actin filaments grow nearly normal to the membrane, realize our model via activity-driven permeation flows Cagnetta et al.
+(2018, ).
+Similarly, in vitro supported bilayers with actin nucleators exhibit ATP-driven fluctuations and forces Waechtler et al.
+(2024) are applicable to our model.
+Our results suggest that coarse-grained experiments, designed to explore long wavelength scaling properties, cannot detect microscopic anisotropy when the surface is stable.
+It would be interesting to explore connections between our predicted unstable phase and observed reduction of membrane tension to even negative values during membrane fusion Solon et al.
 (2006).
 Membrane tension may be measured using optical set ups Popescu et al.
-(2006), giving information about $\nu,\overline{\nu}$.
-Comparison with experimental results on membrane fluctuations by using, e.g., spectroscopy Betz and Sykes (2012) can yield information about $\lambda,\overline{\lambda}$.
+(2006), giving information about $\nu_{x},\nu_{y}$ and $\overline{\nu}$.
+Comparison with experiments on membrane fluctuations by using, e.g., spectroscopy Betz and Sykes (2012) can give $\lambda$.
 Studying how the fluid flows along the membrane in the different phases and their dependence on the tension anisotropy are related to the cortical flows reported in Ref.
 Mayer et al.
-(2010) should be promising future work directions.
-For reasons of simplicity, we set $\nu_{xx}=\nu_{yy}$ and $\lambda_{xx}=\lambda_{yy}$.
-We note that $\nu_{xx}\neq\nu_{yy}$ together with $\lambda_{xx}\neq\lambda_{yy}$ (but without $\overline{\nu},\overline{\lambda}$) was considered in Refs.
-Wolf (1991); Altman et al.
-(2015), giving QLRO translational order and 2D KPZ-like instability for different choices of an effective anisotropy parameter.
-We thus speculate that even in the general case with $\nu_{xx}\neq\nu_{yy}$, $\lambda_{xx}\neq\lambda_{yy}$ and with non-vanishing $\overline{\nu},\overline{\lambda}$ can still have just the three phases studied here.It will be interesting to connect Eq.
+(2010) and how the membrane dynamics may couple with a diffusing density living on the membrane Haldar et al.
+(2023a, b) should be promising future work directions.
+It will be interesting to connect Eq.
 (1) to variants of the Toom model Barabási et al.
@@
 thanks the AvH Stiftung (Germany) for partial financial support under the Research Group Linkage Programme scheme (2024).
-Appendix A Linear instability analysis
+Supplemental Material
+I Linear instability analysis
 We start from Eq.
 (1) of the main text for an anisotropic membrane with a fixed background, ignoring the conservation of embedding bulk fluid momentum.
-As argued, the hydrodynamic equation governing the undulation modes of such a membrane, described by the height field $h({\bf x},t)$, is an anisotropic Kardar-Parisi-Zhang (KPZ) equation with a Gaussian-distributed, zero-mean white noise $f({\bf x},t)$
-$\frac{\partial h}{\partial t}+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (11)
+As argued in the main text, the hydrodynamic equation governing the undulation modes of such a membrane, described by the height field $h({\bf x},t)$, is an anisotropic Kardar-Parisi-Zhang (KPZ) equation with a Gaussian-distributed, zero-mean white noise $f({\bf x},t)$
+$\frac{\partial h}{\partial t}+\frac{\lambda}{2}{({\bm{\nabla}}h)^{2}}=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (14)
 Here, both $i$ and $j$ can take value $1$ or $2$.
 To perform linear stability analysis, we consider only the linear terms in Eq.
-(11), i.e., set $\lambda_{ij}=0$.
+(1), i.e., set $\lambda=0$.
 The resulting linearized equation of $h$ in Fourier space is
-$\partial_{t}h=-\nu(k_{x}^{2}+k_{y}^{2}+2\alpha k_{x}k_{y})h,$ (12)
-[…]
-When $(1+\alpha\sin 2\theta)<0$, linear instability ensues.
-The growth rate of initial fluctuations is maximized when $(1+\alpha\sin 2\theta)$ attains its largest negative value.
+$\partial_{t}h=-\nu_{x}(k_{x}^{2}+\beta k_{y}^{2}+2\alpha k_{x}k_{y})h,$ (15)
+where ${\bf k}=(k_{x},k_{y})$ is a wavevector.
+We define $\beta\equiv\nu_{y}/\nu_{x}$, $\alpha\equiv{\overline{\nu}}/{\nu_{x}}$; see also the main text.
+To proceed further, we use the polar coordinates and introduce $k_{x}=k\cos\theta$ and $k_{y}=k\sin\theta$, where $k=|{\bf k}|$ and $\theta=\tan^{-1}(k_{y}/k_{x})$ is the polar angle.
+Now for linear stability, $\nu_{x}k^{2}(\cos^{2}\theta+\beta\sin^{2}\theta+\alpha\sin 2\theta)<0$ for all $\theta$.
+With $\nu_{x},\,\nu_{y}>0$; as assumed in the main text, we must have $(1+\beta\tan^{2}\theta+2\alpha\tan\theta)>0$ for all $\theta$ to ensure linear stability.
+This means $\beta\geq\alpha^{2}$.
+When $1+\beta\tan^{2}\theta+2\alpha\tan\theta<0$, linear instability occurs.
+The growth rate of the fluctuations is maximized when $1+\beta\tan^{2}\theta+2\alpha\tan\theta$ reaches its most negative value, which occurs at a specific $\theta$ determined by the values of $\alpha$ and $\beta$.
+Consequently, a stripe-like pattern emerges based on the value of $\theta$.
+A representative snapshot of this case is shown in Fig.
+1(c) of the main text.
+Now we consider the special case $\nu_{x}=\nu_{y}=\nu$, i.e.
+$\beta=1$ and $\alpha=\overline{\nu}/\nu$.
+For linear stability, $\nu k^{2}(1+\alpha\sin 2\theta)<0$ for all $\theta$.
+With $\nu>0$, we must have $(1+\alpha\sin 2\theta)>0$ for all $\theta$ to ensure linear stability.
+Since the maximum of $|\sin\,2\theta|$ is unity, the previous condition in turn means $|\alpha|\leq 1$, or $\nu\pm\overline{\nu}>0$.
+When $(1+\alpha\sin 2\theta)<0$, linear instability again ensues.
+The growth rate of the fluctuations is maximized when $(1+\alpha\sin 2\theta)$ attains its largest negative value.
 For $\alpha>1$, this happens at $\theta=3\pi/4$, corresponding to $k_{x}=-k_{y}$ giving the Fourier wavevector with the maximum growth rate.
@@
 $\phi(x,y)=\frac{k}{\sqrt{2}}(y-x).$ (17)
-If we move in the $xy$-plane along paths with a constant phase, i.e., with constant $y-x$, the exponential factor $e^{i\phi}$ does not change.
+If we move in the $xy$-plane along paths with a constant phase, i.e., with constant yxy-x, the exponential factor $e^{i\phi}$ does not change.
 Consequently, the height function $h(x,y)$ does not vary along these directions.
@@
 A representative snapshot of this pattern can be seen in Fig.
-2(a) of the main text.
+1(d) of the main text.
 On the other hand, for $\alpha<-1$, the maximum growth rate occurs at $\theta=\pi/4$, corresponding to $k_{x}=k_{y}$.
-Now using the logic used above, we find this corresponds to the formation of stripes at an angle of $43\pi/$ 4 in the $x$-$y$ plane.
-A representative snapshot of this case is shown in Fig.
-2(b) of the main text.
-Appendix B Galilean invariance
+Now using the logic used above, we can say this corresponds to the formation of stripes at an angle of $3\pi/4$ in the $x$-$y$ plane.
+We have verified this result (not shown).
+II Galilean invariance
 We now explicitly show that under the Galilean transformation defined by
-$t\rightarrow t^{\prime}\equiv t,\,x_{i}\rightarrow x_{i}^{\prime}\equiv x_{i}+\lambda_{ij}u_{j}t,\,h({\bf x},t)\rightarrow h^{\prime}({\bf x}^{\prime},t^{\prime})\equiv h({\bf x},t)+u_{i}x_{i}+\frac{\lambda_{ij}}{2}u_{i}u_{j}t,$ (15)
+$t\rightarrow t^{\prime}\equiv t,\,x_{i}\rightarrow x_{i}^{\prime}\equiv x_{i}+\lambda u_{i}t,\,h({\bf x},t)\rightarrow h^{\prime}({\bf x}^{\prime},t^{\prime})\equiv h({\bf x},t)+u_{i}x_{i}+\frac{\lambda}{2}u^{2}t,$ (18)
 Eq. (1) remains invariant. Under the Galilean transformation
-$\frac{\partial}{\partial t}\rightarrow\frac{\partial}{\partial t^{\prime}}+\lambda_{ij}u_{j}\frac{\partial}{\partial x_{i}^{\prime}},\,\frac{\partial}{\partial x_{i}}\rightarrow\frac{\partial}{\partial x_{i}^{\prime}},$
-$\frac{\partial h({\bf x},t)}{\partial t}\rightarrow\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial t^{\prime}}+\lambda_{ij}u_{j}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x_{i}^{\prime}}-\frac{\lambda_{ij}}{2}u_{i}u_{j}.$ (16)
-$\frac{\lambda_{ij}}{2}\frac{\partial h({\bf x},t)}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{j}}\rightarrow\frac{\lambda_{ij}}{2}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{j}}+\frac{\lambda_{ij}}{2}u_{i}u_{j}-\lambda_{ij}u_{j}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}.$ (17)
+$\frac{\partial}{\partial t}\rightarrow\frac{\partial}{\partial t^{\prime}}+\lambda u_{i}\frac{\partial}{\partial x_{i}^{\prime}},\,\frac{\partial}{\partial x_{i}}\rightarrow\frac{\partial}{\partial x_{i}^{\prime}},$
+$\frac{\partial h({\bf x},t)}{\partial t}\rightarrow\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial t^{\prime}}+\lambda u_{i}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x_{i}^{\prime}}-\frac{\lambda}{2}u^{2}.$ (19)
+$\frac{\lambda}{2}\frac{\partial h({\bf x},t)}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{i}}\rightarrow\frac{\lambda}{2}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}+\frac{\lambda}{2}u^{2}-\lambda u_{i}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}.$ (20)
 $\nu_{ij}\frac{\partial}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{j}}\rightarrow\nu_{ij}\frac{\partial}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{j}}.$ (21)
@@
 (1) is Galilean invariant.
-Appendix C Renormalization group calculations
+III Renormalization group calculations
 We start by constructing a generating functional Bausch et al. (1976); Täuber (2014) by using Eq. (1) and Eq. (2) of the main text. We get
@@
 where $\hat{h}$ is referred as response field and $\mathcal{S}$ is action functional, which is given by
-$S=-\int_{x,t}\hat{h}D\hat{h}+\int_{x,t}\hat{h}(\partial_{t}h+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h-\nu_{ij}\partial_{i}\partial_{j}h).$ (20)
-C.1 Results from the linearized equation of motion
-This can be obtained from the Gaussian part of the action functional (20), i.e., by setting $\lambda_{ij}=0$ in (20).
+$S=-\int_{x,t}\hat{h}D\hat{h}+\int_{x,t}\hat{h}(\partial_{t}h+\frac{\lambda}{2}{({\bm{\nabla}}h)^{2}}-\nu_{ij}\partial_{i}\partial_{j}h).$ (23)
+III.1 Results from the linearized equation of motion
+This can be obtained from the Gaussian part of the action functional (23), i.e., by setting $\lambda=0$ in (23).
 We first define Fourier transforms in space and time by
@@
 Figure 4: Diagrammatic representations of two point functions.
-Figure 5 shows the diagrammatic representations of the various two-point functions.
-C.2 One-loop corrections to the model parameters
+Fig. 4 shows the diagrammatic representations of the various two-point functions.
+III.2 One-loop corrections to the model parameters
 In the (more usual) isotropic models, applications of the Wilson momentum space RG involves integrating the wavevector magnitude $q$ within a shell $\Lambda/b\leq q\leq\Lambda,\,b>1$ Hohenberg and Halperin (1977); Täuber (2014).
@@
 Thus, using spherical coordinates and a spherical shell for mode elimination is not convenient for the present problem.
-Instead, we integrate $q_{x}$ over the whole range from $\infty$- ∞ to $\infty$, and then integrate $q_{y}$ over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$ with $b>1$; see Fig.
-6.
-Of course, we could have chosen another integration scheme, in which $q_{y}$ is integrated over the whole range from $\infty$- ∞ to $\infty$, and then integrate $q_{x}$ over two thin strips $\Lambda/b\leq q_{x}\leq\Lambda$ and $-\Lambda\leq q_{x}\leq-\Lambda/b$.
+Instead, we integrate $q_{x}$ over the whole range from -$\infty$ to $\infty$, and then integrate $q_{y}$ over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $\Lambda\leq q_{y}\leq-\Lambda/b$ with $b>1$; see Fig.
+5.
+Of course, we could have chosen another integration scheme, in which $q_{y}$ is integrated over the whole range from -$\infty$ to $\infty$, and then integrate $q_{x}$ over two thin strips $\Lambda/b\leq q_{x}\leq\Lambda$ and $\Lambda\leq q_{x}\leq-\Lambda/b$.
 For our model, the two schemes are equivalent, connected to the fact that the different anisotropic terms of the same order scale in the same way.
@@
 Figure 5: Wavevector-shell integration in the one-loop diagrams obtained by perturbative expansions of the action functional (23), corresponding to the anisotropic KPZ equation.
-Here, $q_{x}$ is integrated over the range $\infty$- ∞ to $\infty$, while $q_{y}$ is integrated over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$.
+Here, $q_{x}$ is integrated over the range -$\infty$ to $\infty$, while $q_{y}$ is integrated over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $\Lambda\leq q_{y}\leq-\Lambda/b$.
 We now proceed with the evaluation of the integrals that arise from one-loop Feynman diagrams. These integrals take the form
-$\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{m}}=\frac{1}{\nu^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}^{2}+2\alpha q_{x}q_{y}+q_{y}^{2})^{m}}$
-$=\frac{1}{\nu^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}-a_{+})^{m}(q_{x}-a_{-})^{m}},$ (22)
-where $\tilde{f}(q_{x},q_{y})$ represents functions of $q_{x}$ and $q_{y}$, and we use $\alpha={\overline{\nu}}/{\nu}$.
+$\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{m}}=\frac{1}{\nu_{x}^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}^{2}+2\alpha q_{x}q_{y}+\beta q_{y}^{2})^{m}}$
+$=\frac{1}{\nu_{x}^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}-a_{+})^{m}(q_{x}-a_{-})^{m}},$ (25)
+where $\tilde{f}(q_{x},q_{y})$ represents functions of $q_{x}$ and $q_{y}$, and we use $\alpha={\overline{\nu}}/{\nu}$, $\beta=\nu_{y}/\nu_{x}$.
 The exponent $m$ takes values $2$ or $3$.
-The terms $a_{+}=q_{y}(-\alpha+ip)$ and $a_{-}=q_{y}(-\alpha-ip)$ are two poles of order $m$, where $p=\sqrt{1-\alpha^{2}}$.
-To evaluate this integral, we first perform the integration over $q_{x}$ in the range (,)($\infty$,$\infty$)( - ∞ , ∞ ).
-Subsequently, we integrate over the $y$-component of momentum within the two high-$|q_{y}|$ strips, specifically in the regions $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$.
+The terms $a_{+}=q_{y}(-\alpha+ip)$ and $a_{-}=q_{y}(-\alpha-ip)$ are two poles of order $m$, where $p=\sqrt{\beta-\alpha^{2}}$.
+To evaluate this integral, we first perform the integration over $q_{x}$ in the range $(-\infty,\infty)$.
+Subsequently, we integrate over the $y$-component of momentum within the two high-$|q_{y}|$ strips, specifically in the regions $\Lambda/b\leq q_{y}\leq\Lambda$ and $\Lambda\leq q_{y}\leq-\Lambda/b$.
 See Fig.
-6.
+5.
 For computational convenience, we define $q_{x}^{\prime}=q_{x}+\alpha q_{y}$.
@@
 (25) transforms into
-$\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}+\int_{-\Lambda}^{-\Lambda/b}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}.$ (23)
+$\frac{1}{\nu_{x}^{m}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}+\frac{1}{\nu_{x}^{m}}\int_{-\Lambda}^{-\Lambda/b}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}.$ (26)
 The first integral receives contributions from the positive values of $q_{y}$, while the second integral accounts for the negative values of $q_{y}$.
@@
 Hence, in the following analysis, we evaluate the integrals considering only the positive values of $q_{y}$, noting that the contributions from negative $q_{y}$ are identical.
-Below we now evaluate the one-loop diagrams that account for fluctuation-induced corrections to the model parameters.
-C.3 One-loop corrections to $D$
+Below, we list the integrals arising in the calculations and evaluate the one-loop diagrams that account for fluctuation-induced corrections to the model parameters.
+Here, $\int_{q_{x},q_{y}}\equiv\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}}{2\pi}$.
+$I_{1}=\int_{q_{x},q_{y}}\frac{q_{x}^{2}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}}=\frac{1}{4\nu_{x}^{2}}\frac{\beta}{p^{3}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{2}=\int_{q_{x},q_{y}}\frac{q_{y}^{2}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}}=\frac{1}{4\nu_{x}^{2}}\frac{1}{p^{3}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{3}=\int_{q_{x},q_{y}}\frac{q_{x}q_{y}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}}=-\frac{1}{4\nu_{x}^{2}}\frac{\alpha}{p^{3}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{4}=\int_{q_{x},q_{y}}\frac{q_{x}^{4}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=\frac{3}{16\nu_{x}^{3}}\frac{\beta^{2}}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{5}=\int_{q_{x},q_{y}}\frac{q_{x}^{2}q_{y}^{2}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=\frac{3}{16\nu_{x}^{3}}\frac{\beta+2\alpha^{2}}{3p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{6}=\int_{q_{x},q_{y}}\frac{q_{x}^{3}q_{y}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=-\frac{3}{16\nu_{x}^{3}}\frac{\alpha\beta}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{7}=\int_{q_{x},q_{y}}\frac{q_{x}q_{y}^{3}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=-\frac{3}{16\nu_{x}^{3}}\frac{\alpha}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
+$I_{8}=\int_{q_{x},q_{y}}\frac{q_{y}^{4}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=\frac{3}{16\nu_{x}^{3}}\frac{1}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$
+III.3 One-loop corrections to $D$
 The one-loop diagram correcting $D$ is given in Fig. 6, which comes with a symmetry factor of $2$.
@@
 It is given by
-$\lambda_{ij}\lambda_{mn}\int\frac{d\Omega}{2\pi}\frac{d^{2}{q}}{(2\pi)^{2}}(q_{i})(k-q)_{j}\,q_{m}\,(k-q)_{n}\frac{D^{2}}{\Big{[}\Omega^{2}+(\nu_{lk}q_{l}q_{k})^{2}\Big{]}\Big{[}(\omega-\Omega)^{2}+\big{(}\nu_{ps}(k-q)_{p}(k-q)_{s}\big{)}^{2}\Big{]}}.$ (24)
-Setting $k=0$, $\omega=0$, the above integral reduces to
-$\lambda_{ij}\lambda_{mn}\int_{q,\Omega}\frac{q_{i}q_{j}q_{m}q_{n}D^{2}}{\Big{[}\Omega^{2}+(\nu_{xx}q_{x}^{2}+2\nu_{xy}q_{x}q_{y}+\nu_{yy}q_{y}^{2})^{2}\Big{]}^{2}}.$ (25)
-After performing the $\Omega$-integral, the contribution in (25) becomes
-$\frac{\lambda_{ij}\lambda_{mn}}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{D^{2}q_{i}q_{j}q_{m}q_{n}}{(\nu_{xx}q_{x}^{2}+2\nu_{xy}q_{x}q_{y}+\nu_{yy}q_{y}^{2})^{3}}.$ (26)
-Using $\lambda_{xx}=\lambda_{yy}=\lambda$ and $\lambda_{xy}=\lambda_{yx}=\overline{\lambda}$ we get
-$\lambda_{ij}\lambda_{mn}q_{i}q_{j}q_{m}q_{n}=\lambda^{2}q_{x}^{4}+\lambda^{2}q_{y}^{4}+4\overline{\lambda}^{2}q_{x}^{2}q_{y}^{2}+4\lambda\overline{\lambda}q_{x}^{3}q_{y}+4\lambda\overline{\lambda}q_{y}^{3}q_{x}+2\lambda^{2}q_{x}^{2}q_{y}^{2}.$
-Substituting this in Eq. (26) we get the correction to $D$ as
-$\frac{D^{2}}{4\nu^{3}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}}{2\pi}\Biggl{[}\frac{\lambda^{2}q_{x}^{4}+(2\lambda^{2}+4\overline{\lambda}^{2})q_{x}^{2}q_{y}^{2}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}+\frac{4\lambda\overline{\lambda}q_{x}^{3}q_{y}+4\lambda\overline{\lambda}q_{x}q_{y}^{3}+\lambda^{2}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}\Biggl{]},$ (27)
-where $a_{+}=q_{y}[-\alpha+ip]$ and $a_{-}=q_{y}[-\alpha-ip]$ are two poles of third order.
-We set $\nu_{xx}=\nu_{yy}=\nu$, $\nu_{xy}=\nu_{yx}=\overline{\nu}$, and $p=\sqrt{1-\alpha^{2}}$.
-Also defining
-$\mathcal{A}=\lambda^{2},\mathcal{B}=(2\lambda^{2}+4\overline{\lambda}^{2}),\mathcal{C}=4\lambda\overline{\lambda},\mathcal{D}=4\lambda\overline{\lambda},\mathcal{E}=\lambda^{2},$
-Equation (27) becomes
-$\frac{D^{2}}{4\nu^{3}}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}.$ (28)
-Now we perform the $q_{x}$-integral first.
-Considering a closed contour in the upper half complex plane, the relevant poles within the closed contour are $q_{x}=a_{+}=q_{y}(-\alpha+ip)$.
-The residue at $q_{x}=a_{+}$ is
-$\frac{1}{2!}\frac{d^{2}}{dq_{x}^{2}}\biggr{[}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{-})^{3}}\biggr{]}_{q_{x}=a_{+}}.$
-Evaluating, this gives
-$\frac{1}{2}\biggr{[}\frac{12\mathcal{A}q_{x}^{2}+2\mathcal{B}q_{y}^{2}+6\mathcal{C}q_{x}q_{y}}{(q_{x}-a_{-})^{3}}-6\frac{4\mathcal{A}q_{x}^{3}+2\mathcal{B}q_{x}q_{y}^{2}+3\mathcal{C}q_{x}^{2}q_{y}+\mathcal{D}q_{y}^{3}}{(q_{x}-a_{-})^{4}}+12\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{-})^{5}}\biggr{]}_{q_{x}=a_{+}}.$
-Substituting $q_{x}=a_{+}$ in above expression and multiplying with $2\pi i$ gives the result for the $q_{x}$-integral. Thus Eq. (28) becomes
-$\frac{3D^{2}}{64\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\mathcal{A}+\mathcal{E})+\mathcal{B}\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-(\mathcal{C}+\mathcal{D})\alpha\biggr{]}.$ (29)
-After performing the $q_{y}$-integral and using the values of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ Eq. (29) can be written as
-$\frac{3D^{2}}{16\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\frac{\ln b}{2\pi}\biggr{[}\frac{\lambda^{2}}{4}\Bigl{(}2+\frac{2}{3}+\frac{4}{3}\alpha^{2}\Bigl{)}-2\lambda\overline{\lambda}\alpha+\overline{\lambda}^{2}\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}\biggr{]}.$ (30)
-Using $\alpha={\overline{\nu}}/{\nu}$, Eq. (30) becomes
-$\frac{3D^{2}}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\biggr{[}\frac{\lambda^{2}}{4}\Bigl{(}\frac{8\nu^{2}}{3(\nu^{2}-\overline{\nu}^{2})}+\frac{4\overline{\nu}^{2}}{3(\nu^{2}-\overline{\nu}^{2})}\Bigl{)}+\overline{\lambda}^{2}\Bigl{(}\frac{\nu^{2}}{3(\nu^{2}-\overline{\nu}^{2})}+\frac{2\overline{\nu}^{2}}{3(\nu^{2}-\overline{\nu}^{2})}\Bigl{)}-\frac{2\lambda\overline{\lambda}\nu\overline{\nu}}{(\nu^{2}-\overline{\nu}^{2})}\biggr{]}.$ (31)
-C.4 One-loop corrections to $\nu_{ij}$
-The one-loop Feynman diagram, with a symmetry factor 8, that contributes to the fluctuation correction of $\nu$ and $\tilde{\nu}$ is shown in Fig. 8.
-Figure 8: One-loop Feynman diagram that contributes to the correction of $\nu$ and $\overline{\nu}$.
+$\frac{\lambda^{2}}{4}\int\frac{d\Omega}{2\pi}\frac{d^{2}{q}}{(2\pi)^{2}}\frac{4D^{2}\times q^{4}}{\Big[\Omega^{2}+(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{2}\Big]^{2}}.$ (27)
+Where we set $k=0$, $\omega=0$. After performing the $\Omega$-integral, above integral reduces to
+$\frac{\lambda^{2}D^{2}}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q^{4}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}},$ (28)
+$=\frac{\lambda^{2}D^{2}}{4}(I_{4}+2I_{5}+I_{8}),$ (29)
+$=\frac{\lambda^{2}D^{2}}{4}\frac{3}{16\nu_{x}^{3}p^{5}}\frac{\ln b}{2\pi}\Big[\beta^{2}+\frac{2(\beta+2\alpha^{2})}{3}+1\Big].$ (30)
+III.4 One-loop corrections to $\nu_{ij}$
+The one-loop Feynman diagram, with a symmetry factor 8, that contributes to the fluctuation correction of $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$ is shown in Fig. 7.
+Figure 7: One-loop Feynman diagram that contributes to the correction of $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$.
 It is given by
-$-\lambda_{ij}\lambda_{mn}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}q_{m}k_{n}\langle|h_{{\bf q},\Omega}|^{2}\rangle\langle h_{{\bf k-q},-\Omega}\hat{h}_{{\bf-k+q},\Omega}\rangle.$
-After symmetrizing, i.e., ${\bf q}\rightarrow{{\bf q}+{{\bf k}}/{2}}$, and performing the $\Omega$-integral, we are left with two integrals that include corrections for $\nu$ and $\overline{\nu}$.
+$-2D\lambda^{2}\int_{{\bf q},\Omega}\frac{q_{i}(k-q)_{i}q_{j}k_{j}}{\big[i\Omega+\nu_{x}(k-q)_{x}^{2}+\nu_{y}(k-q)_{y}^{2}+2\overline{\nu}(k-q)_{x}(k-q)_{y}\big]\big[\Omega^{2}+(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{2}\big]}.$
+After symmetrizing, i.e., ${\bf q}\rightarrow{{\bf q}+{{\bf k}}/{2}}$, and performing the $\Omega$-integral, we are left with two integrals that include corrections for $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$.
 These integrals are
-$\mathcal{I}_{1}=\frac{\lambda_{ij}\lambda_{mn}k_{n}k_{m}D}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q_{i}q_{j}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}},$ (32)
+$\mathcal{I}_{1}=(k_{x}^{2}+k_{y}^{2})\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}(1+\beta)\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$ (31)
 and
-$\mathcal{I}_{2}=-\frac{\lambda_{ij}\lambda_{mn}k_{n}D}{2}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q_{i}q_{j}q_{m}\bigl{[}\nu(k_{x}q_{x}+k_{y}q_{y})+\overline{\nu}(k_{x}q_{y}+k_{y}q_{x})\bigl{]}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{3}}.$ (33)
-Corrections to $\nu$ are given by the coefficient of $k_{x}^{2}$ (or $k_{y}^{2}$).
-Considering the coefficient of $k_{x}^{2}$, $\lambda_{ij}\lambda_{mn}k_{n}k_{m}q_{i}q_{j}=k_{x}^{2}\lambda(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})$.
-Substituting this in Eq.
-(32), the correction to $\nu$ coming from $\mathcal{I}_{1}$ is
-$\frac{D\lambda k_{x}^{2}}{4}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}}=\frac{D\lambda k_{x}^{2}}{4\nu^{2}}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{(q_{x}-a_{+})^{2}(q_{x}-a_{-})^{2}}.$ (34)
-To evaluate the $q_{x}$-integral, considering a closed contour in the upper half plane the relevant pole $q_{x}=a_{+}=q_{y}(-\alpha+ip)$, which is a second order pole. The residue at $q_{x}=a_{+}$ is
-$\frac{d}{dq_{x}}\Biggl{[}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{(q_{x}-a_{-})^{2}}\Biggl{]}_{q_{x}=a_{+}}=-i\frac{\lambda-\overline{\lambda}\alpha}{2p^{3}q_{y}}.$
-Multiplying by $2\pi i$ and writing $p$ in terms of $\alpha$, Eq. (34) becomes
-$k_{x}^{2}\frac{\lambda D(\lambda-\overline{\lambda}\alpha)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{(2\pi)q_{y}}.$ (35)
-Equation (35) gives the correction to $\nu$ coming from $\mathcal{I}_{1}$. Similarly keeping only order of $k_{x}^{2}$, Eq. (33) becomes
-$-k_{x}^{2}\frac{D}{2\nu^{3}}\int_{q_{x},q_{y}}\biggr{[}\frac{\lambda^{2}\nu q_{x}^{4}+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})q_{x}^{2}q_{y}^{2}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}+\frac{(3\lambda\overline{\lambda}\nu+\lambda^{2}\overline{\nu})q_{x}^{3}q_{y}+(\lambda^{2}\overline{\nu}+2\overline{\lambda}^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)q_{x}q_{y}^{3}+\lambda\overline{\lambda}\overline{\nu}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}\biggr{]}.$ (36)
-This integral is similar to the integral in Eq. (27). However, $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ are now different with values:
-$\mathcal{A}=\lambda^{2}\nu,\mathcal{B}=(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu}),\mathcal{C}=3\lambda\overline{\lambda}\nu+\lambda^{2}\overline{\nu},\mathcal{D}=\lambda^{2}\overline{\nu}+2\overline{\lambda}^{2}\overline{\nu}+\lambda\overline{\lambda}\nu,\mathcal{E}=\lambda\overline{\lambda}\overline{\nu},$
-Comparing with Eq. (29) and substituting these values of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ equation (36) can be written as
-$-k_{x}^{2}\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\Biggl{[}(\lambda^{2}\nu+\lambda\overline{\lambda}\overline{\nu})-(2\lambda^{2}\overline{\nu}+4\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu})\alpha+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})\Bigr{(}\frac{1+2\alpha^{2}}{3}\Bigr{)}\Biggl{]}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (37)
-Equation (37) gives the correction to $\nu$ coming from $\mathcal{I}_{2}$.
-Adding Eqs.
-(35) and (37) and extracting the coefficient of $k_{x}^{2}$, we get the total correction to $\nu$ at the one-loop order as
-$-\frac{3D}{32\nu^{3}{(1-\alpha^{2})^{\frac{5}{2}}}}\frac{\ln b}{2\pi}\Biggl{[}(\lambda^{2}\nu+\lambda\overline{\lambda}\overline{\nu})-(2\lambda^{2}\overline{\nu}+4\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu})\alpha+(\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+3\lambda\overline{\lambda}\overline{\nu})\Bigr{(}\frac{1+2\alpha^{2}}{3}\Bigr{)}\Biggl{]}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{D}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\Biggl{[}\lambda^{2}-\lambda\overline{\lambda}\alpha\Biggl{]}.$ (38)
-Substituting $\alpha={\overline{\nu}}/{\nu}$ in above equation and after some algebra the above expression reduces to
-$\frac{\nu D}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{5}{2}}}\frac{\ln b}{2\pi}\Biggl{[}(\nu^{2}-\overline{\nu}^{2})\Big{(}\frac{\lambda\overline{\lambda}\overline{\nu}}{\nu}-\overline{\lambda}^{2}\Big{)}\Biggl{]}.$ (39)
-To get the corrections to $\overline{\nu}$, we need to find coefficient of $2k_{x}k_{y}$ from Eqs.
-(32) and (33).
-Therefore coonsidering only the $k_{x}k_{y}$-term, $\lambda_{ij}\lambda_{mn}k_{n}k_{m}q_{i}q_{j}\rightarrow 2k_{x}k_{y}\overline{\lambda}\bigl{[}\lambda(q_{x}^{2}+q_{y}^{2})+2\overline{\lambda}q_{x}q_{y}\bigl{]}$.
-Substituting this in Eq.
-(32) we get
-$\frac{D\overline{\lambda}(2k_{x}k_{y})}{4}\int_{q_{x},q_{y}}\frac{\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y}}{\Bigl{[}\nu(q_{x}^{2}+q_{y}^{2})+2\overline{\nu}q_{x}q_{y}\Bigl{]}^{2}}.$ (40)
-The integral in Eq. (40) is same as the integral in Eq. (34). Comparing with Eq. (35) we can write Eq. (40) as
-$(2k_{x}k_{y})\frac{D\overline{\lambda}(\lambda-\overline{\lambda}a)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (41)
-Equation (41) gives the correction to $\overline{\nu}$ that comes from $\mathcal{I}_{1}$. Now,
-$\lambda_{ij}\lambda_{mn}k_{n}q_{i}q_{j}q_{m}=k_{x}(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})(\lambda q_{x}+\overline{\lambda}q_{y})+k_{y}(\lambda q_{x}^{2}+\lambda q_{y}^{2}+2\overline{\lambda}q_{x}q_{y})(\lambda q_{y}+\overline{\lambda}q_{x}).$
-Substituting this in Eq. (33) and extracting the coefficient of $2k_{x}k_{y}$ we get
-$-(2k_{x}k_{y})\frac{D}{2\nu^{3}}\int_{q_{x},q_{y}}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}.$ (42)
-Notice that the integral in Eq.
-(42) has the same form as the integral in Eq.
-(28), but now $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ are given by $\mathcal{A}=\mathcal{E}=\bigr{(}\frac{\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu}{2}\bigr{)},\mathcal{B}=(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu}),\mathcal{C}=\mathcal{D}=(\lambda^{2}\nu+\overline{\lambda}^{2}\nu+2\lambda\overline{\lambda}\overline{\nu})$.
-Thus comparing with Eq.
-(29) we write Eq.
-(42) as
-$-(2k_{x}k_{y})\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)+(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu})\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-2(\lambda^{2}\nu+\overline{\lambda}^{2}\nu+2\lambda\overline{\lambda}\overline{\nu})\alpha\biggr{]}.$ (43)
-Equation (43) gives the correction to $\overline{\nu}$ that comes from $\mathcal{I}_{2}$.
-Thus adding Eq.
-(41) and (43) and extracting the coefficient of $2k_{x}k_{y}$, we get total correction to $\overline{\nu}$ given by
-$-\frac{3D}{32\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\lambda^{2}\overline{\nu}+\lambda\overline{\lambda}\nu)+(3\lambda\overline{\lambda}\nu+2\overline{\lambda}^{2}\overline{\nu}+\lambda^{2}\overline{\nu})\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-(2\lambda^{2}\nu+2\overline{\lambda}^{2}\nu+4\lambda\overline{\lambda}\overline{\nu})\alpha\biggr{]}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{D(\overline{\lambda}\lambda-\overline{\lambda}^{2}\alpha)}{8\nu^{2}(1-\alpha^{2})^{\frac{3}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (44)
-Substituting $\alpha={\overline{\nu}}/{\nu}$ in the above expression and after some algebraic simplification the above contribution becomes
-$\frac{\lambda^{2}\overline{\nu}D}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\Biggl{[}1-\frac{\overline{\lambda}\nu}{\lambda\overline{\nu}}\Biggl{]}.$ (45)
-C.5 Nonrenormalization of $\lambda$ and $\overline{\lambda}$
-Figure 9: One-loop Feynman diagrams giving the corrections to $\lambda$ and $\overline{\lambda}$.
-The left diagram in Fig. 9 has a symmetry factor 24 and includes corrections to both $\lambda$ and $\overline{\lambda}$. We find
-$-\lambda_{ij}\lambda_{mn}\lambda_{rs}D\int_{q,\Omega}\frac{(q_{i}q_{j})(-k_{1m}q_{n})(q_{r}k_{2s})}{(-i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})(i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})\Bigl{(}\Omega^{2}+(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{)}}$
-$=\frac{\lambda_{ij}\lambda_{mn}\lambda_{rs}D}{4}\int_{q}\frac{(q_{i}q_{j})(k_{1m}q_{n})(q_{r}k_{2s})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{3}}.$ (46)
+$\mathcal{I}_{2}=-2D\lambda^{2}\int_{q_{x},q_{y}}\frac{(q_{x}k_{x}+q_{y}k_{y})q^{2}}{4(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}}\big[k_{x}(\nu_{x}q_{x}+\overline{\nu}q_{y})+k_{y}(\nu_{y}q_{y}+\overline{\nu}q_{x})\big].$ (32)
+Corrections to $\nu_{x}$ ($\nu_{y}$) are given by the coefficient of $k_{x}^{2}$ ($k_{y}^{2}$). Considering the coefficient of $k_{x}^{2}$, the correction to $\nu_{x}$ coming from $\mathcal{I}_{2}$ is
+$-k_{x}^{2}\frac{\lambda^{2}D}{2}\int_{q_{x},q_{y}}\frac{\nu_{x}q_{x}^{4}+\nu_{x}q_{x}^{2}q_{y}^{2}+\overline{\nu}q_{x}^{3}q_{y}+\overline{\nu}q_{y}^{3}q_{x}}{(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}},$ (33)
+$=-k_{x}^{2}\frac{\lambda^{2}D}{2}\big[\nu_{x}I_{4}+\nu_{x}I_{5}+\overline{\nu}I_{6}+\overline{\nu}I_{7}\big],$ (34)
+$=-k_{x}^{2}\frac{\lambda^{2}D}{2}\frac{3}{16\nu_{x}^{2}p^{3}}\bigl[\beta+1/3\bigl]\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (35)
+Thus collecting coefficient of $k_{x}^{2}$ from $\mathcal{I}_{1}$ and adding it with above equation we obtain correction to $\nu_{x}$,
+$\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}\frac{\ln b}{2\pi}\bigg[\frac{1}{2}-\frac{\beta}{2}\bigg].$ (36)
+To compute the correction to $\nu_{y}$, we extract the coefficient of $k_{y}^{2}$ from integral $\mathcal{I}_{2}$,
+$-k_{y}^{2}(2D\lambda^{2})\int_{q_{x},q_{y}}\frac{(q_{x}^{2}q_{y}+q_{y}^{3})(\nu_{y}q_{y}+\overline{\nu}q_{x})}{4(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}},$ (37)
+$=-k_{y}^{2}\frac{\lambda^{2}D}{2}\big[\nu_{y}I_{5}+\nu_{y}I_{8}+\overline{\nu}I_{6}+\overline{\nu}I_{7}\big],$ (38)
+$=-k_{y}^{2}\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}\bigl[3/2+\beta/2\bigl]\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$ (39)
+Again collecting coefficient of $k_{y}^{2}$ from $\mathcal{I}_{1}$ and adding it with above equation we obtain correction to $\nu_{y}$,
+$\frac{\lambda^{2}D}{16\nu_{x}^{2}p^{3}}\frac{\ln b}{2\pi}\bigg[\frac{\beta}{2}-\frac{1}{2}\bigg].$ (40)
+Next we collect coefficient of $2k_{x}k_{y}$ from $\mathcal{I}_{2}$ to obtain correction to $\overline{\nu}$ which is,
+$-2k_{x}k_{y}\lambda^{2}D\int_{q_{x},q_{y}}\frac{q^{2}q_{x}(\nu_{y}q_{y}+\overline{\nu}q_{x})+q^{2}q_{y}(\nu_{x}q_{x}+\overline{\nu}q_{y})}{4(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{3}},$ (41)
+$=-(2k_{x}k_{y})\frac{\lambda^{2}D}{4}\big[(\nu_{x}+\nu_{y})(I_{6}+I_{7})+\overline{\nu}(I_{4}+I_{8}+2I_{5})\big].$ (42)
+After some algebra, and collecting coefficient of $2k_{x}k_{y}$ from above equation we obtain the correction to $\overline{\nu}$ which is $\frac{\lambda^{2}D\alpha}{16\nu_{x}^{2}p^{3}}\frac{\ln b}{2\pi}.$
+III.5 Nonrenormalization of $\lambda$
+Figure 8: One-loop Feynman diagrams giving the corrections to $\lambda$.
+The left diagram in Fig. 8 has a symmetry factor 24 and includes corrections to $\lambda$. We find
+$-\lambda^{3}D\int_{{\bf q},\Omega}\frac{q^{2}(-k_{1m}q_{m})(q_{n}k_{2n})}{(-i\Omega+\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})(i\Omega+\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})\Bigl(\Omega^{2}+(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}\Bigl)},$
+$=\frac{\lambda^{3}D}{4}\int_{\bf q}\frac{q^{2}(k_{1m}q_{m})(q_{n}k_{2n})}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}.$ (43)
 Similarly the right diagram in Fig. 8 has a symmetry factor $48$ and contributes
-$-\lambda_{ij}\lambda_{mn}\lambda_{rs}(2D)\int_{q,\Omega}\frac{(q_{i}q_{j})(-k_{1m}q_{n})(-q_{r}k_{2s})}{(-i\Omega+\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{(}\Omega^{2}+(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{2}\Bigl{)}}$
-$=-\frac{\lambda_{ij}\lambda_{mn}\lambda_{rs}D}{4}\int_{q}\frac{(q_{i}q_{j})(k_{1m}q_{n})(q_{r}k_{2s})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{3}}.$ (47)
-Equations (46) and (47), when added, give zero contribution to the correction of $\lambda$ and $\overline{\lambda}$. These are consistent with the Galilean invariance of Eq. (1) in the main text.
-C.6 RG recursion relations
-We integrate out the dynamical fields with wavevectors $k_{y}$ in the range ($\Lambda/b\leq k_{y}\leq\Lambda$ and $-\Lambda\leq k_{y}\leq-\Lambda/b$) in the action functional (20).
-Now we rescale space, time and fields to raise the wavevector cutoff of $k_{y}$ to ±$\pm\Lambda$.
+$-\lambda^{3}(2D)\int_{{\bf q},\Omega}\frac{q^{2}(-k_{1m}q_{m})(-q_{n}k_{2n})}{(-i\Omega+\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}\Bigl[\Omega^{2}+(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}\Bigl]},$
+$=-\frac{\lambda^{3}D}{4}\int_{\bf q}\frac{q^{2}(k_{1m}q_{m})(q_{n}k_{2n})}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}.$ (44)
+Equations (43) and (44), when added, give zero contribution to the correction of $\lambda$. These are consistent with the Galilean invariance of Eq. (1).
+III.6 RG recursion relations
+We integrated out the dynamical fields with wavevectors $k_{y}$ in the range ($\Lambda/b\leq k_{y}\leq\Lambda$ and $\Lambda\leq k_{y}\leq-\Lambda/b$) in the action functional (23).
+Now we rescale space, time and fields to raise the wavevector cutoff of $k_{y}$ to $\pm\Lambda$.
 We rescale ${\bf k}\to{\bf k}^{{}^{\prime}}=b{\bf k}\implies{\bf x}\to{\bf x}^{{}^{\prime}}=\frac{{\bf x}}{b}$ similarly, $\omega\to\omega^{{}^{\prime}}=b^{z}\omega\implies t{{}^{\prime}}=\frac{t}{b^{z}}$.
@@
 $D^{{}^{\prime}}=D^{<}b^{d+3z}\xi^{-2},$
-$\lambda_{ij}^{{}^{\prime}}=\lambda_{ij}^{<}b^{-d-2}\xi.$
+$\lambda^{{}^{\prime}}=\lambda^{<}b^{-d-2}\xi.$
 Also, $h^{<}({\bf x},t)$ scale as $b^{-d-z}\xi h({\bf x}^{{}^{\prime}},t^{{}^{\prime}})$. Defining $\xi_{R}=b^{-d-z}\xi=b^{\chi}$ we get model parameters rescale as
 $D^{{}^{\prime}}=D^{<}b^{z-d-2\chi},$
-$\nu^{{}^{\prime}}=\nu^{<}b^{z-2},$
+$\nu_{x}^{{}^{\prime}}=\nu_{x}^{<}b^{z-2},$
+$\nu_{y}^{{}^{\prime}}=\nu_{y}^{<}b^{z-2},$
 $\overline{\nu}^{{}^{\prime}}=\overline{\nu}^{<}b^{z-2},$
-$\lambda^{{}^{\prime}}=\lambda^{<}b^{\chi+z-2},$
-$\overline{\lambda}^{{}^{\prime}}=\overline{\lambda}^{<}b^{\chi+z-2}.$
-Here, $\chi$ is the roughness exponent of an anisotropic KPZ surface.
-Since there is no correction to $\lambda$, $\overline{\lambda}$ so $\lambda^{<}=\lambda$ and $\overline{\lambda}^{<}=\overline{\lambda}$.
-We further get
-$D^{<}$ $=$ $D\biggr{[}1+\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{[}\Bigl{(}\frac{2\nu^{2}+\overline{\nu}^{2}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}+\frac{\overline{\lambda}^{2}}{\lambda^{2}}\Bigl{(}\frac{\nu^{2}+2\overline{\nu}^{2}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}-\frac{6\overline{\lambda}}{\lambda}\Bigl{(}\frac{\nu\overline{\nu}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}\Bigr{]}\ln b\biggr{]},$ (48)
-$\nu^{<}$ $=$ $\nu\biggr{[}1-\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{(}\frac{\overline{\lambda}\overline{\nu}}{\lambda\nu}-(\frac{\overline{\lambda}}{\lambda})^{2}\Bigr{)}\ln b\biggr{]},$ (49)
-$\overline{\nu}^{<}$ $=$ $\overline{\nu}\biggr{[}1-\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{(}1-\frac{\overline{\lambda}\nu}{\lambda\overline{\nu}}\Bigr{)}\ln b\biggr{]}.$ (50)
-Now set $b=e^{\delta l}\approx 1+\delta l$ for small $\delta l$, and define a dimensionless coupling constant by $g=\frac{\lambda^{2}D}{8\pi(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}$ and dimensionless ratios $\alpha=\frac{\overline{\nu}}{\nu}$ and $\beta=\frac{\overline{\lambda}}{\lambda}$ (see main text).
+$\lambda^{{}^{\prime}}=\lambda^{<}b^{\chi+z-2}.$
+Here, $\chi$ is the roughness exponent of an anisotropic KPZ surface. Since there is no correction to $\lambda$, so $\lambda^{<}=\lambda$. We further get
+$D^{<}$ $=$ $D\biggr[1+\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr(\frac{3\beta^{2}}{4}+\frac{1}{2}(\beta+2\alpha^{2})+\frac{3}{4}\Bigr)\frac{\ln b}{(\beta-\alpha^{2})}\biggr],$ (45)
+$\nu_{x}^{<}$ $=$ $\nu_{x}\biggr[1-\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr(\frac{1}{2}-\frac{\beta}{2}\Bigr)\ln b\biggr],$ (46)
+$\nu_{y}^{<}$ $=$ $\nu_{y}\biggr[1-\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr(\frac{1}{2}-\frac{1}{2\beta}\Bigr)\ln b\biggr],$ (47)
+$\overline{\nu}^{<}$ $=$ $\overline{\nu}\biggr[1-\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\ln b\biggr].$ (48)
+The extra factor of $2$ is multiplied in numerator of the above renormalized parameters to include contribution from negative values of $q_{y}$ i.e.
+($\Lambda\leq q_{y}\leq-\Lambda/b$) in the loop integrals.
+See the lower strip in Fig.
+5.
+Now set $b=e^{\delta l}\approx 1+\delta l$ for small $\delta l$, and define a dimensionless coupling constant by $g=\frac{\lambda^{2}D}{16\pi(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}$ and dimensionless ratios $\alpha=\frac{\overline{\nu}}{\nu_{x}}$ and $\beta=\frac{\nu_{y}}{\nu_{x}}$ (see main text).
 These allow us to obtain the differential RG recursion relations reported in the main text.
-Using the flow equations of $g$ and $\alpha$ it turns out that $g$ is stable if $\beta^{2}>1$ and in this range of $\beta$, $\alpha^{*}=1/\beta$ is locus of stable fixed points (stable fixed line).
-Using $d=2$ and the linear theory results $z=2$, $\chi=0$ we find that $\nu_{l}\sim\nu_{0}l^{A}$, $\overline{\nu}_{l}\sim\overline{\nu}_{0}l^{B}$ and $D_{l}\sim D_{0}l^{C}$.
-Where,
-$A[\beta,\alpha^{*}]=\frac{\beta^{2}-\alpha^{*}\beta}{2s},\;B[\beta,\alpha^{*}]=\frac{(\beta/\alpha^{*})-1}{2s},\;C[\beta,\alpha^{*}]=\frac{1}{2s}\biggr{[}\frac{(2+\alpha^{*2})+\beta^{2}(1+2\alpha^{*2})-6\beta\alpha^{*}}{1-\alpha^{*2}}\biggr{]}.$
-Here, $s=\beta^{2}-1$.
-It turns out that at $\alpha=\alpha^{*}=\frac{1}{\beta}$, $A=B=C=\frac{1}{2s}(\beta^{2}-1)=\frac{1}{2}$ implying $\nu_{l}\sim\nu_{0}\sqrt{l}$, $\overline{\nu}_{l}\sim\overline{\nu}_{0}\sqrt{l}$ and $D_{l}\sim D_{0}\sqrt{l}$.
-Appendix D Direct numerical simulations
+IV Direct numerical simulations
 Following Ref.
@@
 (1).
-Grid points are labeled by ${\bf n}$, and the basis vector is denoted by ${\bf e}_{i}$, where $i=1,2$, giving the discretized form of Eq.
+The grid points are labeled by $\bf n$}, and the basis vector is denoted by ${\bf e}_{i}$, where $i=1,2$, giving the discretized form of Eq.
 (1)
-$h_{\bf n}(t+\Delta t)$ $=h_{\bf n}(t)+\frac{\Delta t}{(a_{0})^{2}}\sum_{i=1}^{2}\biggl{\{}\nu\Bigl{[}h_{\bf n+e_{i}}(t)-2h_{\bf n}(t)+h_{\bf n-e_{i}}(t)\Bigl{]}-\frac{\lambda}{8}\Bigl{[}h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl{]}^{2}\biggl{\}}+\frac{2\overline{\nu}\Delta t}{4(a_{0})^{2}}\Bigl{[}h_{\bf n+e_{1}+e_{2}}(t)$
-$-h_{\bf n+e_{1}-e_{2}}(t)+h_{\bf n-e_{1}-e_{2}}(t)-h_{\bf n-e_{1}+e_{2}}(t)\Bigl{]}-\frac{\overline{\lambda}\Delta t}{4(a_{0})^{2}}\prod_{i=1}^{2}\Bigl{[}h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl{]}+\sigma\sqrt{12\Delta t}R_{n}(t).$ (51)
-Here, $\sigma^{2}=\frac{2D}{(a_{0})^{2}}$.
-The variable $R$ corresponds to uniformly distributed random numbers ranging from $-\frac{1}{2}$ to $\frac{1}{2}$.
-The prefactor $\sigma\sqrt{12\Delta t}$ ensures that the noise possesses an equivalent second moment as Gaussian noise integrated over a time interval $\Delta t$.
-Equation (51) can be conveniently expressed using dimensionless quantities $\tilde{h}=\frac{h}{h_{0}}$, $\tilde{t}=\frac{t}{t_{0}}$, and $\tilde{a}_{0}=\frac{a_{0}}{x_{0}}$, where the natural units are defined as $h_{0}=\frac{\nu}{\lambda}$, $t_{0}=\frac{\nu^{2}}{\sigma^{2}\lambda^{2}}$ and $x_{0}=\sqrt{\frac{\nu^{3}}{\sigma^{2}\lambda^{2}}}$.
-In these units Eq.
-(51) becomes
-$\tilde{h}_{\bf n}(\tilde{t}+\Delta\tilde{t})$ $=\tilde{h}_{\bf n}(\tilde{t})+\frac{\Delta\tilde{t}}{(\tilde{a}_{0})^{2}}\sum_{i=1}^{2}\biggl{\{}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-2\tilde{h}_{\bf n}(\tilde{t})+\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}-\frac{1}{8}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}^{2}\biggl{\}}+\frac{\alpha\Delta\tilde{t}}{2(\tilde{a}_{0})^{2}}\Bigl{[}\tilde{h}_{\bf n+e_{1}+e_{2}}(\tilde{t})$
-$-\tilde{h}_{\bf n+e_{1}-e_{2}}(\tilde{t})+\tilde{h}_{\bf n-e_{1}-e_{2}}(\tilde{t})-\tilde{h}_{\bf n-e_{1}+e_{2}}(\tilde{t})\Bigl{]}-\frac{\beta\Delta\tilde{t}}{4(\tilde{a}_{0})^{2}}\prod_{i=1}^{2}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}+\sqrt{12\Delta\tilde{t}}R_{n}(\tilde{t}).$ (52)
-Figure 10: The variation of width ${\mathcal{W}}$ with time $t$ is presented on a log-log scale for different system sizes with two parameter sets: (a) $\alpha=0.50$, $\beta=2$, corresponding to the logarithmic rough region from RG analysis, with system sizes $L=100$, $300$, $500$, $800$, $1000$, and $1600$; (b) $\alpha=0.50$, $\beta=0.042$, corresponding to the algebraic rough region from our RG analysis, with system sizes $L=100$, $200$, $300$, $400$, and $500$.
-The red dashed line indicates a linear fit in the linear region on a log-log scale, yielding a slope of $\tilde{\beta}=0.234\pm 1.644\times 10^{-5}$ (see text).
-For our simulations we chose $a_{0}=1$ and $\sigma=0.10$.
-First we discuss the numerical results for the logarithmic rough phase.
-For that, we set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$ and $\overline{\lambda}=4$ such that $|\alpha|<1$ and $|\beta|>1$, which corresponds to the logarithmic rough phase as predicted by our RG calculations (see main text).
-For a system of size $L$, we reached the saturation region (after the time after which width $\cal W$ saturates) by simulating Eq.
-(52) up to time $T\approx L^{2}$ in time step of $\Delta t=0.05$ as shown in Fig.
-10(a) for different system sizes.
-Width $\cal W$ at a time $t$ was calculated using ${\cal W}=\sqrt{\langle\tilde{h}({\bf x},t)^{2}\rangle-\langle\tilde{h}({\bf x},t)\rangle^{2}}$.
-Figure 10(a) shows the variation of $\cal W$ with time $t$ in the growth and saturation region on a log-log scale and correspond to a single noise realization.
-The plot clearly shows that growth nearly saturates before reaching time $T\approx L^{2}$.
-To further investigate the saturation region, we extended the simulation for an additional $10^{6}$ time steps.
-During these $10^{6}$ time steps, we performed $N$ noise averages, with different choice of $N$ for different system sizes.
-For each of these time steps width ${\cal W}$ was calculated and averaged over $N$.
-The final saturated width ${\mathcal{W}_{\text{sat}}}$ was obtained by performing an average over $100000$ realizations of ${\cal W}_{\text{sat}}$ obtained after saturation for different system sizes $L=100$, $300$, $500$, $800$, $1000$ and $1600$.
-For $L=100$, $300$ noise average $N$ was set to $100$.
-For $L=500$ $N$ was set to $50$.
-And for $L=800$, $1000$ and $1600$ $N$ was set to $20$.
-Variation of ${\cal W}_{\text{sat}}$ with $\ln L$ in the saturation region is shown in the main text Fig.
-3(a).
-Next we discuss the numerical results for the algebraic rough phase.
-In this region, width ${\cal W}$ is expected to scale as $\mathcal{W}(L,t)\sim L^{\chi}f(\frac{t}{L^{z}})$ where, $f(x)$ is a dimensionless scaling function.
-With parameters $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$ and $\overline{\lambda}=1$ such that $|\alpha|<1$ and $|\beta|<1$, (which should correspond to the algebraic rough phase from RG analysis), we run the simulation for $L=100$, $200$, $300$, $400$ and $500$.
-For both of the growth and saturation region ${\cal W}$ was calculated at each time step $\Delta t=0.01$ with $N$ noise averages up to time $T$.
+$h_{\bf n}(t+\Delta t)$ $=h_{\bf n}(t)+\frac{\Delta t}{(a_{0})^{2}}\sum_{i=1}^{2}\biggl\{\nu_{i}\Bigl[h_{\bf n+e_{i}}(t)-2h_{\bf n}(t)+h_{\bf n-e_{i}}(t)\Bigl]-\frac{\lambda}{8}\Bigl[h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl]^{2}\biggl\}+\frac{\overline{\nu}\Delta t}{2(a_{0})^{2}}\Bigl[h_{\bf n+e_{1}+e_{2}}(t)$
+$-h_{\bf n+e_{1}-e_{2}}(t)+h_{\bf n-e_{1}-e_{2}}(t)-h_{\bf n-e_{1}+e_{2}}(t)\Bigl]+\sqrt{\frac{24D\Delta t}{(a_{0})^{2}}}R_{\bf n}(t).$ (49)
+Here $\nu_{1}(\nu_{2})\equiv\nu_{x}(\nu_{y})$.
+The variable $R$ corresponds to uniformly distributed random numbers ranging from $\frac{1}{2}$ to $\frac{1}{2}$.
+The prefactor $\sqrt{\frac{24D\Delta t}{(a_{0})^{2}}}$ ensures that the noise possesses an equivalent second moment as Gaussian noise integrated over a time interval $\Delta t$.
+For our simulations we chose $a_{0}=1$.
+In the the speculated algebraic rough phase region, width $\cal W$} is expected to scale as $\mathcal{W}(L,t)\sim L^{\chi}f(\frac{t}{L^{z}})$ where, $f(x)$ is a dimensionless scaling function.
+With different set of parameters we run the simulation for $L=100$, $200$, $300$, $400$ and $500$.
+For both of the growth and saturation region $\cal W$} was calculated at different times with $N$ noise averages up to time $T$.
 For system size $L=100$, $200$, $300$, $400$ and $500$ the Eq.
 (49) was simulated up to time $T=5000$, $10000$, $15000$, $25000$ and $30000$ with time step $\Delta t=0.01$.
-$L=100$ to $400$ was noise averaged with $N=100$.
-For $L=500$, $N=50$ was taken.
+$L=100$, $200$ was noise averaged with $N=500$.
+For $L=300$, $400$ and $500$ $N=200$, $100$ and $50$ was taken.
 Refer to Fig.
-10(b) for the variation of ${\cal W}$ over time on a log-log scale.
-The red dashed line indicates the slope of the growth region obtained through linear fitting, yielding the growth exponent $\tilde{\beta}=0.234\pm 1.644\times 10^{-5}$.
-Next we switch over to the saturation region of ${\cal W}$.
+3 in the main text for the variation of $\cal W$} over time on a log-log scale.
 With $T$ defined earlier for different $L$, we reached the saturation region of $\cal W$}.
-See Fig.
-10(b).
 Then we numerically solved the equation of motion (1) for further $100000$ time steps with $\Delta t=0.01$.
@@
 Variation of ${\cal W}_{\text{sat}}$ with $L$ in the saturation region is shown in Fig.
-3(b) of the main text.
+3(d) of the main text.